### v1

# SkyDiffusion: Street-to-Satellite Image Synthesis with Diffusion Models and BEV Paradigm

Abstract
Street-to-satellite image synthesis focuses on generating realistic satellite images from corresponding ground street-view images while maintaining a consistent content layout, similar to looking down from the sky. The significant differences in perspectives create a substantial domain gap between the views, making this cross-view generation task particularly challenging. In this paper, we introduce SkyDiffusion, a novel cross-view generation method for synthesizing satellite images from street-view images, leveraging diffusion models and Bird’s Eye View (BEV) paradigm. First, we design a Curved-BEV method to transform street-view images to the satellite view, reformulating the challenging cross-domain image synthesis task into a conditional generation problem. Curved-BEV also includes a "Multi-to-One" mapping strategy for combining multiple street-view images within the same satellite coverage area, effectively solving the occlusion issues in dense urban scenes. Next, we design a BEV-controlled diffusion model to generate satellite images consistent with the street-view content, which also incorporates a light manipulation module to optimize the lighting condition of the synthesized image using a reference satellite. Experimental results demonstrate that SkyDiffusion outperforms state-of-the-art methods on both suburban (CVUSA & CVACT) and urban (VIGOR-Chicago) cross-view datasets, with an average SSIM increase of 14.5% and a FID reduction of 29.6%, achieving realistic and content-consistent satellite image generation. The code and models of this work will be released at https://opendatalab.github.io/skydiffusion/.
1 Introduction
Cross-view synthesis involves transformations between disparate viewpoints, such as from street-view images to satellite images and vice versa. This process is crucial for applications in autonomous driving [2], 3D reconstruction [40, 7], virtual reality [3, 6], etc. The significant changes in street view and satellite perspectives result in drastically different appearances of objects, leading to a substantial disparity between different image domains. As shown in Fig 1 (a), in cross-view scenarios, building facades observed only in street view and rooftops observed only from satellite view are considered uncorrelated and drastically different features [24, 35, 47]. However, objects such as roads and lawns, which are below the viewpoint, can be observed in different views even if they undergo some deformation, making them cross-view shared correlated information. The key to the challenging task of cross-view generation lies in capturing cross-view relevant information and generating realistic views with consistent content layout.
Figure 1: Illustration of the cross-view image synthesis task. (a) In the cross-view scenario, there are features unique to the street-view image, features exclusive to the satellite image, and shared correlated information observable from both views. (b) Comparison with existing methods using GANs (e.g., CUT [21]) or diffusion models (e.g., Instr-p2p [1]) for street-to-satellite image synthesis.
In the field of cross-view synthesis, converting satellite images to street-view images has received extensive attention [30, 20, 43, 23]. Researchers have leveraged polar transformations [38, 31] or geo-transformation [20] to bridge the gap from the satellite to the street view domain, and then employ Generative Adversarial Networks (GANs) to generate realistic street views. This paper focuses on the less explored yet equally challenging task of synthesizing satellite images from street-view images. Some researchers have used semantic segmentation priors to aid in synthesizing satellite images, achieving notable results [24, 35]. However, the semantic maps are often unavailable in real scenarios. Existing street-to-satellite image generation also lacks transformations such as polar or geo-transformation to fit the gap from the street view to the satellite domain. Moreover, as shown in Fig 1(b), existing methods on street-to-satellite image synthesis, which mostly rely on GANs, often result in lower-quality satellite images with unrealistic textures compared to the ground truth (GT).
Recently, diffusion models have gained attention for their exceptional image synthesis and rich 3D world knowledge derived from large-scale text and 2D image data [9, 45]. Many studies have explored text-guided fine-tuning of diffusion models to top-down perspective from other views [15, 14, 11, 27]. Additionally, novel view synthesis methods have been developed by training with multi-view images and camera poses [19, 29]. Although current novel view synthesis techniques demonstrate satisfactory performance at the object level, their performance in cross-view generation for complex scenes with significant viewpoint differences remains suboptimal. A possible solution is to treat the this task as an image translation task [16, 22, 1]. However, as shown in Fig 1(b), due to the domain gap of different views, the resulting images are often realistic but not actual, with significant discrepancies between the synthesized satellite images and the actual corresponding street view content.
In response to these challenges, this paper introduces SkyDiffusion, the first diffusion-model-based method for synthesizing satellite images from street-view images, similar to looking down from the sky. In SkyDiffusion, we design a curved-BEV method that transforms street-view images into the satellite view to achieve domain alignment between different views. Curved-BEV also includes a "Multi-to-One" BEV mapping method that can optimize BEV perception in densely occluded urban scenarios. Subsequently, we design a BEV-controlled diffusion model to generate satellite images consistent with the street-view content, incorporated with a light manipulation module to address the lack of inherent correlation in lighting conditions between the two views. SkyDiffusion outperforms existing state-of-the-art methods on both suburban (CVUSA and CVACT) and urban (VIGOR-Chicago) cross-view datasets, with an average SSIM increase of 14.5% and a FID reduction of 29.6%, successfully generating realistic and content-consistent satellite images.
2 Related Work
Cross-view image synthesis. Cross-view image synthesis involves generating realistic scenes images with significant viewpoint changes, primarily divided into satellite-to-street [30, 20, 43, 23, 38] and street-to-satellite generation [35, 24, 25]. In this field, synthesizing satellite images from street views is a challenging yet important task. Tang [35] used street-view images and aerial semantic maps for image-to-image translation to synthesize target satellite images, but semantic maps are often unavailable in real scenarios. Regmi [24] attempted to synthesize aerial views from a single street image, but the significant domain differences led to inconsistencies. Additionally, the performance of GANs has also limited the quality of street-to-satellite image synthesis in previous studies.
BEV transformation. BEV transformation is a perspective domain conversion technique similar to polar transformation [31, 38] and geo-transformation [20], widely used in retrieval [4], navigation [17] and localization tasks [34, 41, 32]. Traditional BEV transformation techniques, which rely on camera parameters [28] or depth estimation methods [37, 44], face challenges such as the difficulty in obtaining camera parameters and the high computational cost of depth estimation. Recently, some methods have been proposed to map street-view images to the BEV perspective using geometric relations at the image [41] or feature level [32]. However, such methods are based on the ground plane assumption, where all mapped points are located below the camera, limiting the mapping of upper street scene features and causing information loss. Furthermore, to our knowledge, no studies have yet explored the use of BEV methods for the task of street-to-satellite image generation.
Novel view synthesis based on diffusion models. Diffusion models have shown significant potential in recent studies on novel view synthesis [11, 15, 36, 5, 39]. DreamBooth [27] and AerialDiffusion [14] utilize text guidance to achieve top-down image generation by fine-tuning diffusion models. However, it is challenging to fully encapsulate the complexity of real-world scenes using textual information alone. Zero-1-to-3 [19] and Zero 123++ [29] use multi-view images and camera poses to guide diffusion models in learning object structures and textures, enabling new viewpoint synthesis. However, these methods are limited to object-level tasks with strong viewpoint correlations and struggle to handle the significant viewpoint changes in cross-view generation. Besides, diffusion-based image-to-image translation methods can also generate realistic satellite images [16, 1, 22, 45]. However, the substantial domain gap between street-view images and satellite images presents a significant challenge in generating content-consistent satellite images.
3 Method
As illustrated in Fig. 2, this paper introduces a novel method for synthesizing satellite images from corresponding street-view images. Our cross-view synthesis network, SkyDiffusion, initially applies a curved BEV transformation to the input street-view images, converting the perspective to a top-down view using either one-to-one or multi-to-one mappings (Section 3.1). Subsequently, the BEV-Controlled diffusion model with light manipulation is employed for the controlled synthesis of satellite images (Section 3.2). Following this, we detail our network training process (Section 3.3).
Figure 2: Overview of the proposed SkyDiffusion framework, including the curved BEV transformation and BEV-controlled diffusion model with light manipulation.
3.1 Curved BEV Transformation
One-to-One BEV. Existing BEV methods based on ground plane assumptions ($P(x,y,z=0)$) and geometric relationships achieve the mapping from street view to the BEV plane[32, 41]. However, this assumption fails to map the content above the street-view image and results in noticeable distortion as points move away from the center area. Thus, we propose an improved curved BEV transformation to alleviate this problem and better capture cross-view relevant information.
Figure 3: Schematic of the curved BEV transformation. The upper part (a) details the transformation process, highlighting the two mapping points at the top and bottom of the street view. The lower parts (b) and (c) present the results of One-to-One and Multi-to-One BEV transformations, respectively.
As shown in Fig. 3 (a), we use $P(i,j)$ to denote an index on the BEV space, and $P(x,y,z)$ to illustrate a point in a three-dimensional coordinate system with the camera on the z-axis. We assume that the BEV plane is an upward curved surface, with the height (z-axis) increasing rapidly as the 3D points move away from the center (the central part of Fig 3(a)). The following equation formulates the relationship between $P(x,y,z)$ and $P(i,j)$, where $d_{norm}$ represents the normalized distance from the center, $d_{max}$ is the maximum distance from the center, and $\lambda$ is a scaling factor.
$\begin{cases}x=j-\frac{l}{2},\quad y=\frac{l}{2}-i\\ z=d_{norm}^{4}\times\lambda=\left(\frac{\sqrt{x^{2}+y^{2}}}{d_{max}}\right)^{4}\times\lambda\end{cases}$ (1)
Furthermore, we can convert $P(x,y,z)$ to spherical coordinates $P(\theta,\varphi)$. Based on the equirectangular projection properties of street view images, $P(\theta,\varphi)$ corresponds to the panorama image index $P(u,v)$. This allows us to establish the mapping relationship between $P(x,y,z)$ and $P(u,v)$ from the following equation, where $H$ represents the camera height, and $w$, $h$ represent the width and height of the street image. More details can be found in the supplementary material.
{$u&=\left[\text{arctan2}\left(y,x\right)+\pi\right]\frac{w}{2\pi}\\ v&=\left[\frac{\pi}{2}+\text{arctan2}\left(z-H,\sqrt{x^{2}+y^{2}}\right)\right]\frac{h}{\pi}\end{aligned}\right.$ (2)
Following Eq. 1 and Eq. 2, we achieve the projection from the street-view image to the bird’s eye view without requiring camera calibration parameters or depth estimation methods [28, 37, 44] (Fig. 3 (b)). This allows effective mapping of upper street view features, such as distant roads and buildings. Since the mapping relationship for each transformation is fixed, the computational overhead is minimal.
Multi-to-One BEV. Due to the limitations imposed by dense buildings on street view observations, the BEV sensing range is very limited in urban datasets like VIGOR [46]. To address this, we propose a "Multi-to-one" BEV mapping method that unifies multiple street view BEV mappings results $\textit{BEV}_{\mathrm{cam}_{k}}$ into a single satellite coordinate system based on the relationships of camera position (Fig. 3 (c)). Additionally, we use a simple semantic segmentation network to identify and discard the regions with severe BEV distortion, typically those obstructed by buildings. We select the street view mapping result closest to the capture point for overlapping areas following Eq. 3, where $\textit{BEV}\left(x,y\right)$ represents the unified BEV space corresponding to the satellite image, $x,y$ corresponds to the index position of the BEV plane, $\Delta x_{k}$ and $\Delta y_{k}$ represent the offset of the street-view image from the satellite center, and $x_{\mathrm{cam}_{i}}$ and $y_{\mathrm{cam}_{i}}$ represent the coordinates of camera position in the satellite coordinate system.
$\textit{BEV}\left(x,y\right)=\textit{BEV}_{\mathrm{cam}_{k}}\left(x-\Delta x_{k},y-\Delta y_{k}\right),k=\arg\min_{i}\sqrt{(x-x_{\mathrm{cam}_{i}})^{2}+(y-y_{\mathrm{cam}_{i}})^{2}}$ (3)
This allows us to synthesize large-scale scenes using street-view images that typically contain only limited local information.
3.2 BEV-Controlled Diffusion Model with Light Manipulation
Curved BEV controlled synthesis. Image diffusion models learn to progressively denoise images and generate samples from the training domain. The denoising process can occur in pixel space or in a latent space encoded from training data. Inspired by ControlNet [45], we enhance the control over the generated aerial views by replicating the SD module, enabling the transformed BEV images to effectively guide the synthesis process. Specifically, we first use a small encoder to capture the cross-view associative information of the input BEV image $I_{bev}$ such as road directions and building positions, to achieve BEV image feature embedding $c_{bev}$. Notably, considering the potential feature distortions in BEV images, we employ a spatial attention mechanism to enhance feature capture while suppressing distorted features.
$c_{bev}=\mathcal{E}(I_{bev}).$ (4)
Subsequently, by employing zero convolution layers and replicating the structures and weights of the encoder and middle block from the SD module, the extracted BEV features are injected into another SD module. Guided by the BEV image, the diffusion model considers cross-view associative information, synthesizing aerial views with consistent content.
Light Manipulation. Given the differences in perspectives and timing between satellite and street images, there is no direct correlation in lighting conditions. We employ the DNCM for modeling deterministic color mapping [12]. Initially, we use nDNCM to obtain the "image content information" $I_{c}$ from the preliminary synthesis results $I_{0}$ of the satellite images. Then, we input the lighting and color style $r_{s}$ obtained from the reference aerial view into sDNCM for lighting transfer.
$I_{s}=s\textit{DNCM}(\textit{nDNCM}(I_{0},I_{c}),r_{s}),\text{ where }\textit{DNCM}(I,T)=(I_{h\times w,3}\cdot P_{3,k})\cdot T_{k,k}\cdot Q_{k,3}$ (5)
$I_{c}$ and $r_{s}$ are the content information and color reference information extracted by the encoder $E$, as shown in Fig. 2. The DNCM module internally includes ${P}_{3,k}$ and ${Q}_{k,3}$, which represent the color embedding and recovery matrices, respectively. Here, $k$ represents the embedding dimension. The adaptive color mapping matrix ${T}_{k,k}$ is also obtained through the encoder.
3.3 Network Training
Loss Function. We train the BEV-controlled synthesis of satellite images and the light manipulation module in two separate stages. Given a set of cross-view image data, the image diffusion algorithm gradually adds noise to the satellite image, generating a noisy satellite image $Z_{t}$. It then learns a network $\epsilon_{\theta}$ to predict the noise added to the noisy satellite image $Z_{t}$ based on the given BEV image as the conditional input. Since text cannot fully describe the street view scene, our cross-view image synthesis is essentially an image-conditioned synthesis task, with text not being our optimization target. This helps the model focus on generating satellite images based on the BEV conditions.
$\mathcal{L}=\mathbb{E}_{sat_{0},t,c_{bev},\epsilon\sim\mathcal{N}(0,1)}\left[\left\|\epsilon-\epsilon_{\theta}(sat_{t},t,c_{bev})\right\|_{2}^{2}\right]$ (6)
The loss $\mathcal{L}$ is the learning objective of the diffusion model, where $t$ represents the time step.
Following previous image migration work [13, 12], we apply different LUT filters to the satellite images to simulate the results under various lighting conditions for training the light manipulation module. During training, we use L2 consistency loss $\mathcal{L}_{\text{con}}$ to maintain image content consistency and L1 reconstruction loss $\mathcal{L}_{\text{rec}}$ to learn color style transfer.
Implementation Details. We implement SkyDiffusion using the pre-trained Stable Diffusion v1.5 model [26], with an unlocked diffusion decoder and a classifier-free guidance scale of 9.0 [10]. Final inference sampling uses 50 steps with the DDIM strategy[33]. Training is on eight NVIDIA A100 GPUs with a batch size of 128 for 100 epochs. In the curved BEV transformation, the parameter $\lambda$ is set to 15 and $H$ is set to 1.5. For light manipulation, the reference satellite image is randomly selected from normally toned images in the training set, as shown in Figure 8, and the DNCM mapping matrix parameter $k$ is set to 32 (details in Supplementary Material).
4 Experiments
In this section, we introduce the three datasets utilized and the experimental setup. We then conduct both qualitative and quantitative comparisons of SkyDiffusion with state-of-the-art cross-view synthesis methods. Finally, we perform ablation studies to evaluate the effectiveness of each module.
4.1 Datasets
Suburban dataset (CVUSA & CVACT) are two commonly used cross-view datasets, containing centrally aligned aerial and street-view images from regions in the United States and Canberra, Australia, respectively, in a one-to-one mapping relationship. The data primarily consists of suburban scenes with relatively low-rise buildings and open views. The CVUSA dataset’s training and testing sets are randomly split, while the CVACT dataset is split by region.
Urban dataset (VIGOR-Chicago) is a subset of the VIGOR dataset, containing cross-view image pairs collected from Chicago. In this dataset, the street panorama and satellite images are not centrally aligned, and multiple street-view images cover the same satellite image area, with overlapping regions between different satellite images, forming a many-to-one mapping relationship.
In CVUSA [42], the street-view image is 256 $\times$ 1024 pixels, while in CVACT [18] and VIGOR [46], it is 512 $\times$ 1024, all with north in the central column. Unlike previous methods that synthesize low-resolution satellite images (64 $\times$ 64 or 256 $\times$ 256 pixels), this experiment aims to meet practical needs by generating higher-resolution satellite images (512 $\times$ 512 pixels). The images obtained from street images after the Curved-BEV transformation are the same size as the satellite images.
4.2 Experimental Setup
We compare our approach on three datasets with several state-of-the-art cross-view synthesis methods, including GAN-based methods like X-Seq [24], CUT [21], and CDTE [38], and diffusion models for image translation like img2img-turbo [22] and Instruct pix2pix (Instr-p2p) [1]. We also compare with fine-tuned diffusion models for novel view generation, including DreamBooth [27], Aerial Diff [14], and HawkI [15]. Except for the fine-tuned diffusion models, which use a brief textual description, all other models use the same street-view input for a fair comparison. Following previous work [20, 38, 24], we employ commonly used metrics such as SSIM, PSNR and LPIPS to evaluate the content consistency of the synthesized images, and use FID [8] to assess the realism of images.
4.3 Comparison with State-of-the-Art Methods
On the suburban CVUSA and CVACT datasets, our SkyDiffusion method achieves the best results, with a 28.4% reduction in FID and a 17.4% improvement in SSIM compared to the state-of-the-art, demonstrating its superiority in synthesizing realistic and consistent and satellite images. The visual results in Fig. 4 show that GAN-based cross-view methods generate excessive artifacts and blurriness. While diffusion-based methods such as Aerial Diff [14] and Instr-p2p [1] produce highly realistic aerial views, they lack content correlation with the street-view images. In contrast, our method benefits from BEV perspective transformation, enabling effective capture of associative information. SkyDiffusion generates realistic images with consistent content layout, including road direction, tree placement, and building distribution.
Table 1: Quantitative comparison of different methods on CVUSA [42] and CVACT [18].
Method CVUSA CVACT FID ($\downarrow$) SSIM ($\uparrow$) PSNR ($\uparrow$) LPIPS ($\downarrow$) FID ($\downarrow$) SSIM ($\uparrow$) PSNR ($\uparrow$) LPIPS ($\downarrow$) X-Seq [24] 161.16 0.084 11.97 0.706 190.12 0.042 12.41 0.661 CDTE [38] 122.84 0.143 10.01 0.694 160.81 0.091 12.59 0.663 CUT [21] 72.52 0.122 12.12 0.695 62.34 0.101 13.02 0.662 Aerial Diff [14] 136.18 0.103 10.06 0.855 127.29 0.108 10.24 0.878 DreamBooth [27] 102.99 0.103 11.31 0.812 88.12 0.107 12.81 0.791 HawkI [15] 90.98 0.114 11.09 0.783 135.75 0.112 10.97 0.764 img2img-turbo [22] 77.95 0.103 12.64 0.685 73.24 0.103 12.19 0.679 Instr-p2p [1] 38.01 0.138 12.29 0.697 49.62 0.116 12.51 0.682 Ours 29.01 0.167 13.07 0.653 35.51 0.123 13.58 0.657
Figure 4: Qualitative comparison of synthesis results for suburban areas. Rows 1-3 show the results on the CVUSA dataset, and rows 4-6 show the results on the CVACT dataset.
Figure 5: Qualitative comparison of synthesis results for urban areas (VIGOR-Chicago).
Table 2: Quantitative comparison of different methods on VIGOR-Chicago.
Method Chicago FID ($\downarrow$) SSIM ($\uparrow$) PSNR ($\uparrow$) LPIPS ($\downarrow$) CUT [21] 69.63 0.171 13.82 0.666 Aerial Diff [14] 123.16 0.141 11.49 0.831 DreamBooth[27] 132.16 0.116 9.70 0.754 HawkI [15] 123.47 0.124 9.67 0.752 img2img-turbo [22] 80.10 0.135 11.28 0.690 Instr-p2p [1] 50.30 0.163 10.62 0.689 Ours(One-to-One) 46.35 0.185 10.48 0.665 Ours(Multi-to-One) 31.90 0.205 11.15 0.651
Table 3: Quantitative comparison of cross-dataset generalization test.
Method Cross-dataset FID ($\downarrow$) SSIM ($\uparrow$) PSNR ($\uparrow$) LPIPS ($\downarrow$) CVACT $\rightarrow$ VIGOR-Chicago CUT [21] 146.72 0.130 8.55 0.727 Ours (w/o light) 89.76 0.155 9.45 0.739 Ours 63.04 0.162 11.24 0.723 VIGOR-Chicago $\rightarrow$ CVACT CUT [21] 132.87 0.107 8.22 0.737 Ours (w/o light) 107.58 0.116 8.76 0.751 Ours 100.58 0.121 11.36 0.714
In the urban VIGOR-Chicago dataset, SkyDiffusion achieved excellent performance with a 36.6% reduction in FID and a 20.5% improvement in SSIM compared to the state-of-the-art method. The visual results in Fig. 5 show that our method maintains consistent associative information such as crosswalks and lanes. While the non-associative information on building rooftops cannot be consistently generated, it still produces reasonable and realistic results.
Figure 6: Qualitative comparison of synthesis results for cross-dataset generalization.
4.4 Cross-Dataset Generalization
We trained the model on VIGOR [46] and tested it on the CVACT [18], and vice versa, to evaluate cross-dataset generation capability. Figure 6 shows the cross-dataset generation results of SkyDiffusion. The aim of this task is not to generate identical satellite images but to create scenes with consistent content layouts while exhibiting distinctly different styles. The visualization of the experimental results demonstrates that our method effectively preserves scene content such as road directions and intersections. Additionally, the images processed with the light manipulation module maintain the original content while better matching the illumination conditions of the target domain. Notably, we achieve this by using only a single random image from the training set of the target dataset to obtain the lighting conditions. However, due to the completely different architectural styles of the cities in the CVACT and VIGOR datasets, and because the models are trained on a single dataset, the generalization ability requires further improvement.
Table 4: Ablation study of different modules on CVACT. "Street" represents directly using street-view image inputs, "Curved-BEV" denotes using Curved-BEV transformation, and "Light" stands for the Light Manipulation module.
Method Options CVACT Street Curved-BEV Light FID ($\downarrow$) SSIM ($\uparrow$) PSNR ($\uparrow$) LPIPS ($\downarrow$) CUT [21] 62.34 0.101 13.02 0.662 CUT [21] 51.71 0.121 13.14 0.658 Aerial Diff [14] 127.29 0.108 10.24 0.878 Aerial Diff [14] 83.34 0.114 11.15 0.831 Ours 62.21 0.115 11.95 0.682 Ours 63.60 0.118 13.49 0.681 Ours 36.48 0.117 13.19 0.642 Ours 35.51 0.123 13.58 0.657
Figure 7: Qualitative comparison of synthesis results for One-to-one and Multi-to-one BEV.
Figure 8: Qualitative analysis of ablation results for light manipulation.
4.5 Ablation Study
Curved-BEV controlled synthesis. We conducted ablation experiments on the CVACT dataset, with results shown in Table 4. By comparing the performance of different methods using Curved-BEV and street view inputs, we demonstrate that Curved-BEV consistently enhances performance across various architectures. Additionally, the proposed diffusion method achieves the best results. Further comparisons between existing BEV and Curved-BEV can be found in the supplementary materials.
Furthermore, the Multi-to-one method significantly improves metrics compared to the one-to-one mapping, demonstrating its effectiveness in dense urban scenes (Table 3). As shown in the visual results in Figure 7, using multiple street-view images for multi-to-one BEV mapping effectively addresses occlusions from tall objects like trees and buildings. This method effectively enhances the BEV sensing range in urban environments, facilitating the generation of large-scale satellite scenes.
Light manipulation. The ablation experiments in Table 4 and the cross-dataset generation results in Table 3 indicate that the Light Manipulation module aligns the lighting conditions of the synthesized images with those of the target domain images, significantly improving SSIM and PSNR metrics. The visualization results in Figure 8 further support this. This improvement is due to the consistent lighting conditions in the satellite data from CVACT and VIGOR-Chicago, which were captured in the same city. Using a small satellite image slice as a reference effectively captures the lighting conditions at the time of capture. Details on this module can be found in the supplementary materials.
5 Conclusion
In this study, we introduce SkyDiffusion, a novel approach specifically designed for street images to satellite images cross-view synthesis. The network operates solely with street images as input, utilizing a curved BEV method and diffusion models to generate realistic and content-consistent satellite images, similar to looking down from the sky. SkyDiffusion achieves state-of-the-art performance in both content consistency and image realism on suburban (CVUSA & CVACT) and urban (VIGOR-Chicago) cross-view datasets, demonstrating significant potential in real-world image synthesis in complex application scenes.

### v2

@@
 First, we design a Curved-BEV method to transform street-view images to the satellite view, reformulating the challenging cross-domain image synthesis task into a conditional generation problem.
-Curved-BEV also includes a "Multi-to-One" mapping strategy for combining multiple street-view images within the same satellite coverage area, effectively solving the occlusion issues in dense urban scenes.
-Next, we design a BEV-controlled diffusion model to generate satellite images consistent with the street-view content, which also incorporates a light manipulation module to optimize the lighting condition of the synthesized image using a reference satellite.
-Experimental results demonstrate that SkyDiffusion outperforms state-of-the-art methods on both suburban (CVUSA & CVACT) and urban (VIGOR-Chicago) cross-view datasets, with an average SSIM increase of 14.5% and a FID reduction of 29.6%, achieving realistic and content-consistent satellite image generation.
-The code and models of this work will be released at https://opendatalab.github.io/skydiffusion/.
-1 Introduction
+Curved-BEV also includes a "Multi-to-One" mapping strategy for leveraging multiple street-view images within the same satellite coverage area, effectively solving the occlusion issues in dense urban scenes.
+Next, we design a BEV-controlled diffusion model to generate satellite images consistent with the street-view content, which also incorporates a light manipulation module to make the lighting conditions of the synthesized satellite images more flexible.
+Experimental results demonstrate that SkyDiffusion outperforms state-of-the-art methods on both suburban (CVUSA & CVACT) and urban (VIGOR-Chicago) cross-view datasets, with an average SSIM increase of 13.96% and a FID reduction of 20.54%, achieving realistic and content-consistent satellite image generation.
+The code and models of this work will be released at https://opendatalab.github.io/skydiffusion/
+Introduction
 Cross-view synthesis involves transformations between disparate viewpoints, such as from street-view images to satellite images and vice versa.
-This process is crucial for applications in autonomous driving [2], 3D reconstruction [40, 7], virtual reality [3, 6], etc.
+This process is crucial for applications in urban planning (Li et al.
+2023b), 3D reconstruction (Wang et al.
+2021), virtual reality (Emmaneel et al.
+2023), etc.
 The significant changes in street view and satellite perspectives result in drastically different appearances of objects, leading to a substantial disparity between different image domains.
-As shown in Fig 1 (a), in cross-view scenarios, building facades observed only in street view and rooftops observed only from satellite view are considered uncorrelated and drastically different features [24, 35, 47].
+As shown in Fig 1 (a), in cross-view scenarios, building facades observed only in street view and rooftops observed only from satellite view are considered uncorrelated and drastically different features (Regmi and Borji 2018; Zhu et al.
+2023).
 However, objects such as roads and lawns, which are below the viewpoint, can be observed in different views even if they undergo some deformation, making them cross-view shared correlated information.
-The key to the challenging task of cross-view generation lies in capturing cross-view relevant information and generating realistic views with consistent content layout.
+The key to the challenging task of cross-view synthesis lies in bridging domain gaps, capturing cross-view relevant information, and generating realistic views with consistent content layouts.
 Figure 1: Illustration of the cross-view image synthesis task.
 (a) In the cross-view scenario, there are features unique to the street-view image, features exclusive to the satellite image, and shared correlated information observable from both views.
-(b) Comparison with existing methods using GANs (e.g., CUT [21]) or diffusion models (e.g., Instr-p2p [1]) for street-to-satellite image synthesis.
-In the field of cross-view synthesis, converting satellite images to street-view images has received extensive attention [30, 20, 43, 23].
-Researchers have leveraged polar transformations [38, 31] or geo-transformation [20] to bridge the gap from the satellite to the street view domain, and then employ Generative Adversarial Networks (GANs) to generate realistic street views.
+(b) Comparison of our street-to-satellite image synthesis method with existing cross-view synthesis methods (e.g., SelectionGAN (Tang et al.
+2019)) or image translation methods based on diffusion models (e.g., Instr-p2p (Brooks, Holynski, and Efros 2023)).
+In the field of cross-view synthesis, converting satellite images to street-view images has received extensive attention (Shi et al.
+2022; Lu et al.
+2020; Qian et al.
+2023).
+Researchers have leveraged polar transformations (Toker et al.
+2021) or geo-transformation (Lu et al.
+2020) to bridge the gap from the satellite to the street view domain, and then employ Generative Adversarial Networks (GANs) to generate realistic street views.
 This paper focuses on the less explored yet equally challenging task of synthesizing satellite images from street-view images.
-Some researchers have used semantic segmentation priors to aid in synthesizing satellite images, achieving notable results [24, 35].
+Some researchers have used semantic segmentation priors to aid in synthesizing satellite images, achieving notable results (Tang et al.
+2019).
 However, the semantic maps are often unavailable in real scenarios.
 Existing street-to-satellite image generation also lacks transformations such as polar or geo-transformation to fit the gap from the street view to the satellite domain.
-Moreover, as shown in Fig 1(b), existing methods on street-to-satellite image synthesis, which mostly rely on GANs, often result in lower-quality satellite images with unrealistic textures compared to the ground truth (GT).
-Recently, diffusion models have gained attention for their exceptional image synthesis and rich 3D world knowledge derived from large-scale text and 2D image data [9, 45].
-Many studies have explored text-guided fine-tuning of diffusion models to top-down perspective from other views [15, 14, 11, 27].
-Additionally, novel view synthesis methods have been developed by training with multi-view images and camera poses [19, 29].
+Moreover, as shown in Fig 1(b), existing methods on street-to-satellite image synthesis, which mostly rely on GANs, often result in lower-quality satellite images with unrealistic textures.
+Recently, diffusion models have gained attention for their exceptional image synthesis (Ho, Jain, and Abbeel 2020; Zhang, Rao, and Agrawala 2023).
+Many studies have explored text-guided fine-tuning of diffusion models to top-down perspective from other views (Kothandaraman et al.
+2023a; Ruiz et al.
+2023).
+Additionally, novel view synthesis methods have been developed by training with multi-view images and camera poses (Liu et al.
+2023; Shi et al.
+2023a).
 Although current novel view synthesis techniques demonstrate satisfactory performance at the object level, their performance in cross-view generation for complex scenes with significant viewpoint differences remains suboptimal.
-A possible solution is to treat the this task as an image translation task [16, 22, 1].
+A possible solution is to treat the this task as an image translation task (Li et al.
+2023a; Parmar et al.
+2024).
 However, as shown in Fig 1(b), due to the domain gap of different views, the resulting images are often realistic but not actual, with significant discrepancies between the synthesized satellite images and the actual corresponding street view content.
-In response to these challenges, this paper introduces SkyDiffusion, the first diffusion-model-based method for synthesizing satellite images from street-view images, similar to looking down from the sky.
-In SkyDiffusion, we design a curved-BEV method that transforms street-view images into the satellite view to achieve domain alignment between different views.
-Curved-BEV also includes a "Multi-to-One" BEV mapping method that can optimize BEV perception in densely occluded urban scenarios.
-Subsequently, we design a BEV-controlled diffusion model to generate satellite images consistent with the street-view content, incorporated with a light manipulation module to address the lack of inherent correlation in lighting conditions between the two views.
-SkyDiffusion outperforms existing state-of-the-art methods on both suburban (CVUSA and CVACT) and urban (VIGOR-Chicago) cross-view datasets, with an average SSIM increase of 14.5% and a FID reduction of 29.6%, successfully generating realistic and content-consistent satellite images.
-2 Related Work
-Cross-view image synthesis.
-Cross-view image synthesis involves generating realistic scenes images with significant viewpoint changes, primarily divided into satellite-to-street [30, 20, 43, 23, 38] and street-to-satellite generation [35, 24, 25].
+In response to these challenges, this paper introduces SkyDiffusion, a street-to-satellite image synthesis method based on diffusion models and the BEV paradigm, similar to viewing from the sky.
+Unlike previous cross-view synthesis methods, our approach is based on a diffusion framework, enabling the generation of more realistic satellite images.
+We design a Curved-BEV method that transforms street-view images into satellite views, achieving domain alignment between different perspectives and facilitating content-consistent satellite image synthesis.
+To address the limited perception range of BEV in densely occluded urban scenarios, the Curved-BEV method also includes a Multi-to-One mapping strategy.
+Additionally, our BEV-controlled diffusion model incorporates a light manipulation module, making the lighting conditions of the synthesized satellite images more adaptable.
+Our main contributions are summarized as follows:
+  • We introduce SkyDiffusion, a novel street-to-satellite image synthesis method leveraging diffusion models and BEV paradigm to generate realistic and content-consistent satellite images.
+  • We design a Curved-BEV method to transform street-view images into satellite views for domain alignment.
+It also includes "Multi-to-One" mapping strategy to enhance BEV perception range in densely occluded urban areas.
+  • SkyDiffusion outperforms state-of-the-art methods on suburban (CVUSA, CVACT) and urban (VIGOR-Chicago) cross-view datasets, with a 13.96% SSIM increase and 20.54% FID reduction.
+Related Work
+Cross-view image synthesis involves generating realistic scenes images with significant viewpoint changes, primarily divided into satellite-to-street (Shi et al.
+2022; Lu et al.
+2020; Qian et al.
+2023) and street-to-satellite generation (Regmi and Borji 2018, 2019).
 In this field, synthesizing satellite images from street views is a challenging yet important task.
-Tang [35] used street-view images and aerial semantic maps for image-to-image translation to synthesize target satellite images, but semantic maps are often unavailable in real scenarios.
-Regmi [24] attempted to synthesize aerial views from a single street image, but the significant domain differences led to inconsistencies.
+Regmi and Borji (2018) attempted to synthesize aerial views from a single street image, but the significant domain differences led to inconsistencies.
+Tang et al.
+(2019) used street-view images and aerial semantic maps for image-to-image translation to synthesize target satellite images, but semantic maps are often unavailable in real scenarios.
 Additionally, the performance of GANs has also limited the quality of street-to-satellite image synthesis in previous studies.
-BEV transformation.
-BEV transformation is a perspective domain conversion technique similar to polar transformation [31, 38] and geo-transformation [20], widely used in retrieval [4], navigation [17] and localization tasks [34, 41, 32].
-Traditional BEV transformation techniques, which rely on camera parameters [28] or depth estimation methods [37, 44], face challenges such as the difficulty in obtaining camera parameters and the high computational cost of depth estimation.
-Recently, some methods have been proposed to map street-view images to the BEV perspective using geometric relations at the image [41] or feature level [32].
+Figure 2: Overview of the proposed SkyDiffusion framework, including the curved BEV transformation and BEV-controlled diffusion model with light manipulation.
+BEV transformation is a perspective domain conversion technique similar to polar transformation (Shi et al.
+2019; Toker et al.
+2021) and geo-transformation (Lu et al.
+2020), widely used in retrieval (Ye et al.
+2024b), navigation (Li et al.
+2022) and localization tasks (Song et al.
+2023; Wang et al.
+2023; Shi et al.
+2023b).
+Traditional BEV transformation techniques, which rely on camera parameters (Sarlin et al.
+2023) or depth estimation methods (Teng et al.
+2024; Ye et al.
+2024a), face challenges such as the difficulty in obtaining camera parameters and the high computational cost of depth estimation.
+Recently, some methods have been proposed to map street-view images to the BEV perspective using geometric relations at the image (Wang et al.
+2023) or feature level (Shi et al.
+2023b).
 However, such methods are based on the ground plane assumption, where all mapped points are located below the camera, limiting the mapping of upper street scene features and causing information loss.
-Furthermore, to our knowledge, no studies have yet explored the use of BEV methods for the task of street-to-satellite image generation.
+Furthermore, to our knowledge, no studies have yet explored the use of BEV methods for street-to-satellite image synthesis task.
 Novel view synthesis based on diffusion models.
-Diffusion models have shown significant potential in recent studies on novel view synthesis [11, 15, 36, 5, 39].
-DreamBooth [27] and AerialDiffusion [14] utilize text guidance to achieve top-down image generation by fine-tuning diffusion models.
+Diffusion models have shown significant potential in recent studies on novel view synthesis (Kawar et al.
+2023; Gao et al.
+2024; Tseng et al.
+2023).
+DreamBooth (Ruiz et al.
+2023) and AerialDiffusion (Kothandaraman et al.
+2023a) utilize text guidance to achieve top-down image generation by fine-tuning diffusion models.
 However, it is challenging to fully encapsulate the complexity of real-world scenes using textual information alone.
-Zero-1-to-3 [19] and Zero 123++ [29] use multi-view images and camera poses to guide diffusion models in learning object structures and textures, enabling new viewpoint synthesis.
+Zero-1-to-3 (Liu et al.
+2023) and Zero 123++ (Shi et al.
+2023a) use multi-view images and camera poses to guide diffusion models in learning object structures and textures, enabling new viewpoint synthesis.
 However, these methods are limited to object-level tasks with strong viewpoint correlations and struggle to handle the significant viewpoint changes in cross-view generation.
-Besides, diffusion-based image-to-image translation methods can also generate realistic satellite images [16, 1, 22, 45].
+Besides, diffusion-based image-to-image translation methods can also generate realistic satellite images (Li et al.
+2023a; Brooks, Holynski, and Efros 2023; Zhang, Rao, and Agrawala 2023).
 However, the substantial domain gap between street-view images and satellite images presents a significant challenge in generating content-consistent satellite images.
-3 Method
+Methods
 As illustrated in Fig.
-2, this paper introduces a novel method for synthesizing satellite images from corresponding street-view images.
-Our cross-view synthesis network, SkyDiffusion, initially applies a curved BEV transformation to the input street-view images, converting the perspective to a top-down view using either one-to-one or multi-to-one mappings (Section 3.1).
-Subsequently, the BEV-Controlled diffusion model with light manipulation is employed for the controlled synthesis of satellite images (Section 3.2).
-Following this, we detail our network training process (Section 3.3).
-Figure 2: Overview of the proposed SkyDiffusion framework, including the curved BEV transformation and BEV-controlled diffusion model with light manipulation.
-3.1 Curved BEV Transformation
+2, this paper introduces SkyDiffusion, a novel method for synthesizing satellite images from corresponding street-view images.
+It initially applies a curved BEV transformation to the input street-view images, converting the perspective to a top-down view using either one-to-one or multi-to-one mappings.
+Subsequently, the BEV-Controlled diffusion model with light manipulation is employed for the controlled synthesis of satellite images.
+Following this, we detail our network training process.
+Curved BEV Transformation
 One-to-One BEV.
-Existing BEV methods based on ground plane assumptions ($P(x,y,z=0)$) and geometric relationships achieve the mapping from street view to the BEV plane[32, 41].
+Existing BEV methods based on ground plane assumptions ($P(x,y,z=0)$) and geometric relationships achieve the mapping from street view to the BEV plane(Shi et al.
+2023b; Wang et al.
+2023).
 However, this assumption fails to map the content above the street-view image and results in noticeable distortion as points move away from the center area.
-Thus, we propose an improved curved BEV transformation to alleviate this problem and better capture cross-view relevant information.
+Thus, we propose an enhanced curved BEV transformation to better capture cross-view information and address this issue.
 Figure 3: Schematic of the curved BEV transformation.
@@
 This allows us to establish the mapping relationship between $P(x,y,z)$ and $P(u,v)$ from the following equation, where $H$ represents the camera height, and $w$, $h$ represent the width and height of the street image.
-More details can be found in the supplementary material.
+More details are in the supplementary material.
 {$u&=\left[\text{arctan2}\left(y,x\right)+\pi\right]\frac{w}{2\pi}\\ v&=\left[\frac{\pi}{2}+\text{arctan2}\left(z-H,\sqrt{x^{2}+y^{2}}\right)\right]\frac{h}{\pi}\end{aligned}\right.$ (2)
@@
 1 and Eq.
-2, we achieve the projection from the street-view image to the bird’s eye view without requiring camera calibration parameters or depth estimation methods [28, 37, 44] (Fig.
+2, we achieve the projection from the street-view image to the bird’s eye view without requiring camera calibration parameters or depth estimation methods (Teng et al.
+2024; Ye et al.
+2024a) (Fig.
 3 (b)).
-This allows effective mapping of upper street view features, such as distant roads and buildings.
+This effectively maps upper street view features like distant roads and buildings.
 Since the mapping relationship for each transformation is fixed, the computational overhead is minimal.
 Multi-to-One BEV.
-Due to the limitations imposed by dense buildings on street view observations, the BEV sensing range is very limited in urban datasets like VIGOR [46].
+Due to the limitations imposed by dense buildings on street view observations, the BEV sensing range is very limited in urban datasets like VIGOR.
 To address this, we propose a "Multi-to-one" BEV mapping method that unifies multiple street view BEV mappings results $\textit{BEV}_{\mathrm{cam}_{k}}$ into a single satellite coordinate system based on the relationships of camera position (Fig.
 3 (c)).
-Additionally, we use a simple semantic segmentation network to identify and discard the regions with severe BEV distortion, typically those obstructed by buildings.
+We use a simple semantic segmentation network to identify and discard the regions with severe BEV distortion, typically those obstructed by buildings.
 We select the street view mapping result closest to the capture point for overlapping areas following Eq.
 3, where $\textit{BEV}\left(x,y\right)$ represents the unified BEV space corresponding to the satellite image, $x,y$ corresponds to the index position of the BEV plane, $\Delta x_{k}$ and $\Delta y_{k}$ represent the offset of the street-view image from the satellite center, and $x_{\mathrm{cam}_{i}}$ and $y_{\mathrm{cam}_{i}}$ represent the coordinates of camera position in the satellite coordinate system.
-$\textit{BEV}\left(x,y\right)=\textit{BEV}_{\mathrm{cam}_{k}}\left(x-\Delta x_{k},y-\Delta y_{k}\right),k=\arg\min_{i}\sqrt{(x-x_{\mathrm{cam}_{i}})^{2}+(y-y_{\mathrm{cam}_{i}})^{2}}$ (3)
-This allows us to synthesize large-scale scenes using street-view images that typically contain only limited local information.
-3.2 BEV-Controlled Diffusion Model with Light Manipulation
+$&\textit{BEV}\left(x,y\right)=\textit{BEV}_{\mathrm{cam}_{k}}\left(x-\Delta x_{k},y-\Delta y_{k}\right)\\ &k=\arg\min_{i}\sqrt{(x-x_{\mathrm{cam}_{i}})^{2}+(y-y_{\mathrm{cam}_{i}})^{2}}\end{aligned}\right.$ (3)
+This enables us to synthesize large-scale scenes using street-view images with limited local information.
+BEV-Controlled Synthesis with Light Manipulation
 Curved BEV controlled synthesis.
-Image diffusion models learn to progressively denoise images and generate samples from the training domain.
+Image diffusion models progressively denoise images and generate samples from the training domain.
 The denoising process can occur in pixel space or in a latent space encoded from training data.
-Inspired by ControlNet [45], we enhance the control over the generated aerial views by replicating the SD module, enabling the transformed BEV images to effectively guide the synthesis process.
+Inspired by ControlNet (Zhang, Rao, and Agrawala 2023), we enhance the control over the generated aerial views by replicating the SD module, enabling the transformed BEV images to effectively guide the synthesis process.
 Specifically, we first use a small encoder to capture the cross-view associative information of the input BEV image $I_{bev}$ such as road directions and building positions, to achieve BEV image feature embedding $c_{bev}$.
@@
 $c_{bev}=\mathcal{E}(I_{bev}).$ (4)
-Subsequently, by employing zero convolution layers and replicating the structures and weights of the encoder and middle block from the SD module, the extracted BEV features are injected into another SD module.
-Guided by the BEV image, the diffusion model considers cross-view associative information, synthesizing aerial views with consistent content.
+Subsequently, by using zero convolution layers and replicating the encoder and middle block structures and weights from the SD module, the extracted BEV features are injected into another SD module.
+Guided by the BEV image, the diffusion model incorporates cross-view associative information, synthesizing aerial views with consistent content.
 Light Manipulation.
 Given the differences in perspectives and timing between satellite and street images, there is no direct correlation in lighting conditions.
-We employ the DNCM for modeling deterministic color mapping [12].
+To address this, we designed a light manipulation module to make the lighting conditions of the synthesized satellite images more adaptable.
+We employ the DNCM for modeling deterministic color mapping (Ke et al.
+2023).
 Initially, we use nDNCM to obtain the "image content information" $I_{c}$ from the preliminary synthesis results $I_{0}$ of the satellite images.
 Then, we input the lighting and color style $r_{s}$ obtained from the reference aerial view into sDNCM for lighting transfer.
-$I_{s}=s\textit{DNCM}(\textit{nDNCM}(I_{0},I_{c}),r_{s}),\text{ where }\textit{DNCM}(I,T)=(I_{h\times w,3}\cdot P_{3,k})\cdot T_{k,k}\cdot Q_{k,3}$ (5)
+{$&I_{s}=s\textit{DNCM}(\textit{nDNCM}(I_{0},I_{c}),r_{s})\\ &\textit{DNCM}(I,T)=(I_{h\times w,3}\cdot P_{3,k})\cdot T_{k,k}\cdot Q_{k,3}\end{aligned}\right.$ (5)
+Figure 4: Qualitative Comparison of Synthesis Results.
+The first and second rows present the results on the CVUSA dataset, while the third and fourth rows display the results on the CVACT dataset.
+The fifth and sixth rows showcase the visualized outcomes on the VIGOR-Chicago dataset.
 $I_{c}$ and $r_{s}$ are the content information and color reference information extracted by the encoder $E$, as shown in Fig.
@@
 The adaptive color mapping matrix ${T}_{k,k}$ is also obtained through the encoder.
-3.3 Network Training
+Notably, the lighting adjustment module is independent of the satellite image synthesis, meaning that satellite images can be generated without relying on this module.
+We obtain lighting information from randomly selected satellite images in the training set, and standard images can also be used as lighting control conditions.
+Network Training
 Loss Function.
@@
 The loss $\mathcal{L}$ is the learning objective of the diffusion model, where $t$ represents the time step.
-Following previous image migration work [13, 12], we apply different LUT filters to the satellite images to simulate the results under various lighting conditions for training the light manipulation module.
+Following previous work (Ke et al.
+2022, 2023), we apply different LUT filters to the satellite images to simulate the results under various lighting conditions for training the light manipulation module.
 During training, we use L2 consistency loss $\mathcal{L}_{\text{con}}$ to maintain image content consistency and L1 reconstruction loss $\mathcal{L}_{\text{rec}}$ to learn color style transfer.
+Method CVUSA CVACT VIGOR-Chicago
+FID ($\downarrow$) SSIM ($\uparrow$) PSNR ($\uparrow$) LPIPS ($\downarrow$) FID ($\downarrow$) SSIM ($\uparrow$) PSNR ($\uparrow$) LPIPS ($\downarrow$) FID ($\downarrow$) SSIM ($\uparrow$) PSNR ($\uparrow$) LPIPS ($\downarrow$)
+X-Seq 161.16 0.084 11.97 0.706 190.12 0.042 12.41 0.661 - - - -
+CDTE 122.84 0.143 10.01 0.694 160.81 0.091 12.59 0.663 - - - -
+SelectionGAN 116.57 0.129 12.38 0.742 100.21 0.116 12.04 0.684 149.53 0.127 11.77 0.778
+Aerial Diff 136.18 0.103 10.06 0.855 127.29 0.108 10.24 0.878 123.16 0.141 11.49 0.831
+HawkI 90.98 0.114 11.09 0.783 135.75 0.112 10.97 0.764 123.47 0.124 9.67 0.752
+CUT 72.83 0.121 12.09 0.687 62.22 0.102 13.11 0.664 69.42 0.169 13.66 0.665
+img2img-turbo 77.95 0.127 12.64 0.685 73.24 0.103 12.19 0.679 80.10 0.135 11.28 0.690
+Instr-p2p 38.01 0.138 12.29 0.697 49.62 0.116 12.51 0.682 50.30 0.163 10.62 0.689
+Ours 29.18 0.168 14.58 0.635 35.51 0.123 13.58 0.657 45.29 0.186 11.69 0.661
+Table 1: Quantitative comparison of different methods on CVUSA, CVACT and VIGOR-Chicago.
 Implementation Details.
-We implement SkyDiffusion using the pre-trained Stable Diffusion v1.5 model [26], with an unlocked diffusion decoder and a classifier-free guidance scale of 9.0 [10].
-Final inference sampling uses 50 steps with the DDIM strategy[33].
+We implement SkyDiffusion using the pre-trained Stable Diffusion v1.5 model (Rombach et al.
+2022), with an unlocked diffusion decoder and a classifier-free guidance scale of 9.0 (Ho and Salimans 2022).
+Final inference sampling uses 50 steps with the DDIM strategy (Song, Meng, and Ermon 2021).
 Training is on eight NVIDIA A100 GPUs with a batch size of 128 for 100 epochs.
-In the curved BEV transformation, the parameter $\lambda$ is set to 15 and $H$ is set to 1.5.
-For light manipulation, the reference satellite image is randomly selected from normally toned images in the training set, as shown in Figure 8, and the DNCM mapping matrix parameter $k$ is set to 32 (details in Supplementary Material).
-4 Experiments
+For light manipulation, the reference satellite image is randomly selected from normally toned images in the training set, and the DNCM mapping matrix parameter $k$ is set to 32 (details in Supplementary Material).
+Experiments
 In this section, we introduce the three datasets utilized and the experimental setup.
@@
 Finally, we perform ablation studies to evaluate the effectiveness of each module.
-4.1 Datasets
-Suburban dataset (CVUSA & CVACT) are two commonly used cross-view datasets, containing centrally aligned aerial and street-view images from regions in the United States and Canberra, Australia, respectively, in a one-to-one mapping relationship.
-The data primarily consists of suburban scenes with relatively low-rise buildings and open views.
+Datasets
+CVUSA & CVACT (Suburban dataset).
+CVUSA (Workman, Souvenir, and Jacobs 2015) and CVACT (Liu and Li 2019) are two widely used cross-view datasets.
+They contain centrally aligned satellite and street-view images from regions in the United States and Canberra, Australia, respectively, with a one-to-one mapping relationship.
+The data primarily consists of suburban scenes characterized by relatively low-rise buildings and open views.
 The CVUSA dataset’s training and testing sets are randomly split, while the CVACT dataset is split by region.
-Urban dataset (VIGOR-Chicago) is a subset of the VIGOR dataset, containing cross-view image pairs collected from Chicago.
-In this dataset, the street panorama and satellite images are not centrally aligned, and multiple street-view images cover the same satellite image area, with overlapping regions between different satellite images, forming a many-to-one mapping relationship.
-In CVUSA [42], the street-view image is 256 $\times$ 1024 pixels, while in CVACT [18] and VIGOR [46], it is 512 $\times$ 1024, all with north in the central column.
+VIGOR-Chicago (Urban dataset).
+VIGOR-Chicago (Zhu, Yang, and Chen 2021) is a subset of the VIGOR dataset, containing cross-view image pairs collected from Chicago urban area.
+In this dataset, the street panorama and satellite images are not centrally aligned, and multiple street-view images cover the same satellite image area, with overlapping regions between different satellite images, forming a multi-to-one mapping relationship.
+All input street images are set to 512 × 1024 pixels, with the center column aligned to the north.
 Unlike previous methods that synthesize low-resolution satellite images (64 $\times$ 64 or 256 $\times$ 256 pixels), this experiment aims to meet practical needs by generating higher-resolution satellite images (512 $\times$ 512 pixels).
-The images obtained from street images after the Curved-BEV transformation are the same size as the satellite images.
-4.2 Experimental Setup
-We compare our approach on three datasets with several state-of-the-art cross-view synthesis methods, including GAN-based methods like X-Seq [24], CUT [21], and CDTE [38], and diffusion models for image translation like img2img-turbo [22] and Instruct pix2pix (Instr-p2p) [1].
-We also compare with fine-tuned diffusion models for novel view generation, including DreamBooth [27], Aerial Diff [14], and HawkI [15].
-Except for the fine-tuned diffusion models, which use a brief textual description, all other models use the same street-view input for a fair comparison.
-Following previous work [20, 38, 24], we employ commonly used metrics such as SSIM, PSNR and LPIPS to evaluate the content consistency of the synthesized images, and use FID [8] to assess the realism of images.
-4.3 Comparison with State-of-the-Art Methods
-On the suburban CVUSA and CVACT datasets, our SkyDiffusion method achieves the best results, with a 28.4% reduction in FID and a 17.4% improvement in SSIM compared to the state-of-the-art, demonstrating its superiority in synthesizing realistic and consistent and satellite images.
+In the Curved-BEV transformation, $\lambda$ is set to 3 and $H$ to 1.5 for CVUSA and CVACT datasets.
+For the VIGOR-Chicago dataset, $\lambda$ is set to 10, with $H$ unchanged.
+The Curved-BEV transformation results in street images that match the size of the satellite images.
+Experimental Setup
+We compared our method on three datasets with several state-of-the-art cross-view synthesis methods, including the commonly used cross-view satellite synthesis methods X-Seq (Regmi and Borji 2018), SelectionGAN (without semantic masks) (Tang et al.
+2019), and CDTE (Toker et al.
+2021), image translation methods such as the GAN-based CUT (Park et al.
+2020) and diffusion-based Instruct pix2pix (Instr-p2p) (Brooks, Holynski, and Efros 2023) and img2img-turbo (Parmar et al.
+2024), as well as novel view synthesis methods like Aerial Diff (Kothandaraman et al.
+2023a) and HawkI (Kothandaraman et al.
+2023b).
+Except for the fine-tuned diffusion model, which requires brief textual descriptions, all other models use the same street-view input for a fair comparison.
+Following previous work (Lu et al.
+2020; Toker et al.
+2021; Regmi and Borji 2018), we employ commonly used metrics such as SSIM, PSNR and LPIPS to evaluate the content consistency of the synthesized images, and use FID (Heusel et al.
+2017) to assess the realism of images.
+Comparison with State-of-the-Art Methods
+On the suburban CVUSA and CVACT datasets, our SkyDiffusion method achieved the outstanding results, as shown in Table 1.
+Compared to state-of-the-art methods, it reduced FID by 25.83% and increased SSIM by 13.89%, demonstrating its superiority in synthesizing realistic and consistent satellite images.
 The visual results in Fig.
 4 show that GAN-based cross-view methods generate excessive artifacts and blurriness.
-While diffusion-based methods such as Aerial Diff [14] and Instr-p2p [1] produce highly realistic aerial views, they lack content correlation with the street-view images.
-In contrast, our method benefits from BEV perspective transformation, enabling effective capture of associative information.
+While diffusion-based methods such as Aerial Diff and Instr-p2p produce highly realistic aerial views, they lack content correlation with the street-view images.
+In contrast, our method benefits from BEV perspective transformation, enabling effective capture of cross-view associative information.
 SkyDiffusion generates realistic images with consistent content layout, including road direction, tree placement, and building distribution.
-Table 1: Quantitative comparison of different methods on CVUSA [42] and CVACT [18].
-Method CVUSA CVACT FID ($\downarrow$) SSIM ($\uparrow$) PSNR ($\uparrow$) LPIPS ($\downarrow$) FID ($\downarrow$) SSIM ($\uparrow$) PSNR ($\uparrow$) LPIPS ($\downarrow$) X-Seq [24] 161.16 0.084 11.97 0.706 190.12 0.042 12.41 0.661 CDTE [38] 122.84 0.143 10.01 0.694 160.81 0.091 12.59 0.663 CUT [21] 72.52 0.122 12.12 0.695 62.34 0.101 13.02 0.662 Aerial Diff [14] 136.18 0.103 10.06 0.855 127.29 0.108 10.24 0.878 DreamBooth [27] 102.99 0.103 11.31 0.812 88.12 0.107 12.81 0.791 HawkI [15] 90.98 0.114 11.09 0.783 135.75 0.112 10.97 0.764 img2img-turbo [22] 77.95 0.103 12.64 0.685 73.24 0.103 12.19 0.679 Instr-p2p [1] 38.01 0.138 12.29 0.697 49.62 0.116 12.51 0.682 Ours 29.01 0.167 13.07 0.653 35.51 0.123 13.58 0.657
-Figure 4: Qualitative comparison of synthesis results for suburban areas. Rows 1-3 show the results on the CVUSA dataset, and rows 4-6 show the results on the CVACT dataset.
-Figure 5: Qualitative comparison of synthesis results for urban areas (VIGOR-Chicago).
-Table 2: Quantitative comparison of different methods on VIGOR-Chicago.
-Method Chicago FID ($\downarrow$) SSIM ($\uparrow$) PSNR ($\uparrow$) LPIPS ($\downarrow$) CUT [21] 69.63 0.171 13.82 0.666 Aerial Diff [14] 123.16 0.141 11.49 0.831 DreamBooth[27] 132.16 0.116 9.70 0.754 HawkI [15] 123.47 0.124 9.67 0.752 img2img-turbo [22] 80.10 0.135 11.28 0.690 Instr-p2p [1] 50.30 0.163 10.62 0.689 Ours(One-to-One) 46.35 0.185 10.48 0.665 Ours(Multi-to-One) 31.90 0.205 11.15 0.651
-Table 3: Quantitative comparison of cross-dataset generalization test.
-Method Cross-dataset FID ($\downarrow$) SSIM ($\uparrow$) PSNR ($\uparrow$) LPIPS ($\downarrow$) CVACT $\rightarrow$ VIGOR-Chicago CUT [21] 146.72 0.130 8.55 0.727 Ours (w/o light) 89.76 0.155 9.45 0.739 Ours 63.04 0.162 11.24 0.723 VIGOR-Chicago $\rightarrow$ CVACT CUT [21] 132.87 0.107 8.22 0.737 Ours (w/o light) 107.58 0.116 8.76 0.751 Ours 100.58 0.121 11.36 0.714
-In the urban VIGOR-Chicago dataset, SkyDiffusion achieved excellent performance with a 36.6% reduction in FID and a 20.5% improvement in SSIM compared to the state-of-the-art method.
+Additional visual results can be found in the supplementary materials.
+In the urban VIGOR-Chicago dataset, SkyDiffusion reduced FID by 9.96% and improved SSIM by 14.11% compared to the state-of-the-art method, as shown in Table 1.
+For fair comparison, the VIGOR results for SkyDiffusion use only a single street-view image.
+The results using Multi-to-One BEV generation with multiple street-view images can be found in the ablation study.
 The visual results in Fig.
@@
 While the non-associative information on building rooftops cannot be consistently generated, it still produces reasonable and realistic results.
-Figure 6: Qualitative comparison of synthesis results for cross-dataset generalization.
-4.4 Cross-Dataset Generalization
-We trained the model on VIGOR [46] and tested it on the CVACT [18], and vice versa, to evaluate cross-dataset generation capability.
-Figure 6 shows the cross-dataset generation results of SkyDiffusion.
-The aim of this task is not to generate identical satellite images but to create scenes with consistent content layouts while exhibiting distinctly different styles.
-The visualization of the experimental results demonstrates that our method effectively preserves scene content such as road directions and intersections.
-Additionally, the images processed with the light manipulation module maintain the original content while better matching the illumination conditions of the target domain.
-Notably, we achieve this by using only a single random image from the training set of the target dataset to obtain the lighting conditions.
-However, due to the completely different architectural styles of the cities in the CVACT and VIGOR datasets, and because the models are trained on a single dataset, the generalization ability requires further improvement.
-Table 4: Ablation study of different modules on CVACT.
-"Street" represents directly using street-view image inputs, "Curved-BEV" denotes using Curved-BEV transformation, and "Light" stands for the Light Manipulation module.
-Method Options CVACT Street Curved-BEV Light FID ($\downarrow$) SSIM ($\uparrow$) PSNR ($\uparrow$) LPIPS ($\downarrow$) CUT [21] 62.34 0.101 13.02 0.662 CUT [21] 51.71 0.121 13.14 0.658 Aerial Diff [14] 127.29 0.108 10.24 0.878 Aerial Diff [14] 83.34 0.114 11.15 0.831 Ours 62.21 0.115 11.95 0.682 Ours 63.60 0.118 13.49 0.681 Ours 36.48 0.117 13.19 0.642 Ours 35.51 0.123 13.58 0.657
-Figure 7: Qualitative comparison of synthesis results for One-to-one and Multi-to-one BEV.
-Figure 8: Qualitative analysis of ablation results for light manipulation.
-4.5 Ablation Study
+Ablation Study
 Curved-BEV controlled synthesis.
-We conducted ablation experiments on the CVACT dataset, with results shown in Table 4.
-By comparing the performance of different methods using Curved-BEV and street view inputs, we demonstrate that Curved-BEV consistently enhances performance across various architectures.
-Additionally, the proposed diffusion method achieves the best results.
-Further comparisons between existing BEV and Curved-BEV can be found in the supplementary materials.
-Furthermore, the Multi-to-one method significantly improves metrics compared to the one-to-one mapping, demonstrating its effectiveness in dense urban scenes (Table 3).
+We conducted ablation experiments on the three datasets, with results shown in Table 2.
+Compared to directly using street-view images as input, the Curved-BEV method improves performance across multiple metrics by transforming street-view images into satellite views for domain alignment.
+This indicates that the Curved-BEV method aids in synthesizing more content-consistent satellite images.
+Furthermore, the Multi-to-one method further improves metrics compared to the one-to-one mapping, demonstrating its effectiveness in dense urban scenes.
 As shown in the visual results in Figure 5, using multiple street-view images for multi-to-one BEV mapping effectively addresses occlusions from tall objects like trees and buildings.
 This method effectively enhances the BEV sensing range in urban environments, facilitating the generation of large-scale satellite scenes.
+Additional ablation study results can be found in the supplementary materials.
+Method FID ($\downarrow$) SSIM ($\uparrow$) PSNR ($\uparrow$) LPIPS ($\downarrow$)
+CVUSA
+Baseline 32.45 0.149 12.63 0.650
+Ours (C-BEV) 29.18 0.168 14.58 0.635
+CVACT
+Baseline 62.21 0.115 11.95 0.682
+Ours (C-BEV) 35.51 0.123 13.58 0.657
+VIGOR-Chicago
+Baseline 53.27 0.170 10.38 0.666
+Ours (C-BEV) 45.29 0.186 11.69 0.661
+Ours (C-BEV Multi) 31.90 0.205 11.15 0.651
+Table 2: Ablation study of the Curved-BEV module. "Baseline" represents directly using street-view image,"C-BEV" denotes using Curved-BEV transformation, and "Multi" stands for Multi-to-One strategy.
+Figure 5: Qualitative comparison of synthesis results for One-to-one and Multi-to-one BEV.
 Light manipulation.
-The ablation experiments in Table 4 and the cross-dataset generation results in Table 3 indicate that the Light Manipulation module aligns the lighting conditions of the synthesized images with those of the target domain images, significantly improving SSIM and PSNR metrics.
-The visualization results in Figure 8 further support this.
-This improvement is due to the consistent lighting conditions in the satellite data from CVACT and VIGOR-Chicago, which were captured in the same city.
+The ablation experiments in Table 3 indicate that the Light Manipulation module aligns the lighting conditions of the synthesized images with those of the target domain images, improving SSIM and PSNR metrics.
+This is because satellites typically capture large areas with relatively uniform overall brightness.
 Using a small satellite image slice as a reference effectively captures the lighting conditions at the time of capture.
-Details on this module can be found in the supplementary materials.
-5 Conclusion
+Due to the dispersed distribution of the CVUSA data and the more complex lighting conditions, the improvements are not as pronounced as on other datasets.
+The light processing module maintains the original image content while providing more flexible lighting conditions for the synthesized satellite images.
+More visualization results can be found in the supplementary material.
+Method FID ($\downarrow$) SSIM ($\uparrow$) PSNR ($\uparrow$) LPIPS ($\downarrow$)
+CVUSA
+Ours (w/o light) 29.02 0.167 14.07 0.643
+Ours 29.18 0.168 14.58 0.635
+CVACT
+Ours (w/o light) 36.48 0.117 12.85 0.665
+Ours 35.51 0.123 13.58 0.657
+VIGOR-Chicago
+Ours (w/o light) 46.77 0.179 10.46 0.666
+Ours 45.29 0.186 11.69 0.661
+Table 3: Ablation study of the light manipulation module.
+Figure 6: Synthesis results for cross-dataset generalization.
+Method FID ($\downarrow$) SSIM ($\uparrow$) PSNR ($\uparrow$) LPIPS ($\downarrow$)
+CVACT $\rightarrow$ VIGOR-Chicago
+Instr-p2p 100.80 0.152 9.14 0.717
+Ours (w/o light) 89.76 0.155 9.45 0.739
+Ours 63.04 0.162 11.24 0.723
+VIGOR-Chicago $\rightarrow$ CVACT
+Instr-p2p 121.29 0.113 8.90 0.759
+Ours (w/o light) 107.58 0.116 8.76 0.751
+Ours 100.58 0.121 11.36 0.714
+Table 4: Cross-dataset generalization assessment.
+Cross-Dataset Generalization
+We trained the model on CVACT and tested it on the VIGOR-Chicago, and vice versa, to evaluate cross-dataset generation capability, as shown in Table 4.
+Compared to Instruct pix2pix, our method (w/o light) demonstrates superior performance across metrics.
+Figure 6 shows the synthesis results for cross-dataset generalization.
+Our method effectively preserves scene content such as road directions and intersections.
+However, due to the distinct city styles of the CVACT and VIGOR datasets and the fact that the model was trained on a single dataset, there are noticeable differences in elements such as buildings and lighting conditions.
+The aim of this task is not to generate identical satellite images but to create scenes with consistent content layouts while exhibiting distinctly different styles.
+Additionally, we can flexibly use the light manipulation module to obtain light information from a random image in the target dataset’s training set, better matching the target domain’s lighting conditions.
+Conclusion
 In this study, we introduce SkyDiffusion, a novel approach specifically designed for street images to satellite images cross-view synthesis.
-The network operates solely with street images as input, utilizing a curved BEV method and diffusion models to generate realistic and content-consistent satellite images, similar to looking down from the sky.
-SkyDiffusion achieves state-of-the-art performance in both content consistency and image realism on suburban (CVUSA & CVACT) and urban (VIGOR-Chicago) cross-view datasets, demonstrating significant potential in real-world image synthesis in complex application scenes.
+SkyDiffusion operates solely with street images as input, utilizing a BEV Paradigm and diffusion models to generate satellite images.
+SkyDiffusion achieves state-of-the-art performance in both content consistency and image realism on suburban and urban cross-view datasets, demonstrating its superior capabilities in cross-view synthesis tasks.
+The satellite and street-view domains bridging method, Curved-BEV, within this approach also holds valuable insights for tasks such as cross-view retrieval and geolocation.
+In the future work, we plan to extend our method to more cities and further refine it for specific downstream applications, such as urban planning and virtual tourism.

### v3

@@
-SkyDiffusion: Street-to-Satellite Image Synthesis with Diffusion Models and BEV Paradigm
+SkyDiffusion: Ground-to-Aerial Image Synthesis with Diffusion Models and BEV Paradigm
 Abstract
-Street-to-satellite image synthesis focuses on generating realistic satellite images from corresponding ground street-view images while maintaining a consistent content layout, similar to looking down from the sky.
-The significant differences in perspectives create a substantial domain gap between the views, making this cross-view generation task particularly challenging.
-In this paper, we introduce SkyDiffusion, a novel cross-view generation method for synthesizing satellite images from street-view images, leveraging diffusion models and Bird’s Eye View (BEV) paradigm.
-First, we design a Curved-BEV method to transform street-view images to the satellite view, reformulating the challenging cross-domain image synthesis task into a conditional generation problem.
-Curved-BEV also includes a "Multi-to-One" mapping strategy for leveraging multiple street-view images within the same satellite coverage area, effectively solving the occlusion issues in dense urban scenes.
-Next, we design a BEV-controlled diffusion model to generate satellite images consistent with the street-view content, which also incorporates a light manipulation module to make the lighting conditions of the synthesized satellite images more flexible.
-Experimental results demonstrate that SkyDiffusion outperforms state-of-the-art methods on both suburban (CVUSA & CVACT) and urban (VIGOR-Chicago) cross-view datasets, with an average SSIM increase of 13.96% and a FID reduction of 20.54%, achieving realistic and content-consistent satellite image generation.
-The code and models of this work will be released at https://opendatalab.github.io/skydiffusion/
-Introduction
-Cross-view synthesis involves transformations between disparate viewpoints, such as from street-view images to satellite images and vice versa.
-This process is crucial for applications in urban planning (Li et al.
-2023b), 3D reconstruction (Wang et al.
-2021), virtual reality (Emmaneel et al.
-2023), etc.
-The significant changes in street view and satellite perspectives result in drastically different appearances of objects, leading to a substantial disparity between different image domains.
-As shown in Fig 1 (a), in cross-view scenarios, building facades observed only in street view and rooftops observed only from satellite view are considered uncorrelated and drastically different features (Regmi and Borji 2018; Zhu et al.
-2023).
-However, objects such as roads and lawns, which are below the viewpoint, can be observed in different views even if they undergo some deformation, making them cross-view shared correlated information.
-The key to the challenging task of cross-view synthesis lies in bridging domain gaps, capturing cross-view relevant information, and generating realistic views with consistent content layouts.
+Ground-to-aerial image synthesis focuses on generating realistic aerial images from corresponding ground street view images while maintaining consistent content layout, simulating a top-down view.
+The significant viewpoint difference leads to domain gaps between views, and dense urban scenes limit the visible range of street views, making this cross-view generation task particularly challenging.
+In this paper, we introduce SkyDiffusion, a novel cross-view generation method for synthesizing aerial images from street view images, utilizing a diffusion model and the Bird’s-Eye View (BEV) paradigm.
+The Curved-BEV method in SkyDiffusion converts street-view images into a BEV perspective, effectively bridging the domain gap, and employs a "multi-to-one" mapping strategy to address occlusion issues in dense urban scenes.
+Next, SkyDiffusion designed a BEV-guided diffusion model to generate content-consistent and realistic aerial images.
+Additionally, we introduce a novel dataset, Ground2Aerial-3, designed for diverse ground-to-aerial image synthesis applications, including disaster scene aerial synthesis, historical high-resolution satellite image synthesis, and low-altitude UAV image synthesis tasks.
+Experimental results demonstrate that SkyDiffusion outperforms state-of-the-art methods on cross-view datasets across natural (CVUSA), suburban (CVACT), urban (VIGOR-Chicago), and various application scenarios (G2A-3), achieving realistic and content-consistent aerial image generation.
+More result and dataset information can be found at https://opendatalab.github.io/skydiffusion/.
+1 Introduction
+In fields such as land cover classification [18, 36], urban planning [47, 45], and disaster response [11, 44, 38], the analysis of aerial imagery plays a crucial role.
+The cross-view Ground-to-Aerial Image Synthesis task leverages ground street-view images to synthesize corresponding aerial images at the same location.
+Street-view images can be captured and uploaded by users on platforms such as Google Street View or Mapillary, offering high flexibility in data collection and frequent temporal coverage [1].
+Therefore, Ground-to-Aerial Image Synthesis can support rapid response during natural disasters [19] or fill historical gaps in high-resolution satellite imagery records [3], among other applications.
 Figure 1: Illustration of the cross-view image synthesis task.
-(a) In the cross-view scenario, there are features unique to the street-view image, features exclusive to the satellite image, and shared correlated information observable from both views.
-(b) Comparison of our street-to-satellite image synthesis method with existing cross-view synthesis methods (e.g., SelectionGAN (Tang et al.
-2019)) or image translation methods based on diffusion models (e.g., Instr-p2p (Brooks, Holynski, and Efros 2023)).
-In the field of cross-view synthesis, converting satellite images to street-view images has received extensive attention (Shi et al.
-2022; Lu et al.
-2020; Qian et al.
-2023).
-Researchers have leveraged polar transformations (Toker et al.
-2021) or geo-transformation (Lu et al.
-2020) to bridge the gap from the satellite to the street view domain, and then employ Generative Adversarial Networks (GANs) to generate realistic street views.
-This paper focuses on the less explored yet equally challenging task of synthesizing satellite images from street-view images.
-Some researchers have used semantic segmentation priors to aid in synthesizing satellite images, achieving notable results (Tang et al.
-2019).
-However, the semantic maps are often unavailable in real scenarios.
-Existing street-to-satellite image generation also lacks transformations such as polar or geo-transformation to fit the gap from the street view to the satellite domain.
-Moreover, as shown in Fig 1(b), existing methods on street-to-satellite image synthesis, which mostly rely on GANs, often result in lower-quality satellite images with unrealistic textures.
-Recently, diffusion models have gained attention for their exceptional image synthesis (Ho, Jain, and Abbeel 2020; Zhang, Rao, and Agrawala 2023).
-Many studies have explored text-guided fine-tuning of diffusion models to top-down perspective from other views (Kothandaraman et al.
-2023a; Ruiz et al.
-2023).
-Additionally, novel view synthesis methods have been developed by training with multi-view images and camera poses (Liu et al.
-2023; Shi et al.
-2023a).
-Although current novel view synthesis techniques demonstrate satisfactory performance at the object level, their performance in cross-view generation for complex scenes with significant viewpoint differences remains suboptimal.
-A possible solution is to treat the this task as an image translation task (Li et al.
-2023a; Parmar et al.
-2024).
-However, as shown in Fig 1(b), due to the domain gap of different views, the resulting images are often realistic but not actual, with significant discrepancies between the synthesized satellite images and the actual corresponding street view content.
-In response to these challenges, this paper introduces SkyDiffusion, a street-to-satellite image synthesis method based on diffusion models and the BEV paradigm, similar to viewing from the sky.
-Unlike previous cross-view synthesis methods, our approach is based on a diffusion framework, enabling the generation of more realistic satellite images.
-We design a Curved-BEV method that transforms street-view images into satellite views, achieving domain alignment between different perspectives and facilitating content-consistent satellite image synthesis.
-To address the limited perception range of BEV in densely occluded urban scenarios, the Curved-BEV method also includes a Multi-to-One mapping strategy.
-Additionally, our BEV-controlled diffusion model incorporates a light manipulation module, making the lighting conditions of the synthesized satellite images more adaptable.
-Our main contributions are summarized as follows:
-  • We introduce SkyDiffusion, a novel street-to-satellite image synthesis method leveraging diffusion models and BEV paradigm to generate realistic and content-consistent satellite images.
+(a) Challenges of cross-view domain gaps; (b) Challenges of occlusion in dense scenes; (c) Comparing our ground-to-aerial image synthesis method with existing cross-view synthesis methods.
+Despite the potential of cross-view synthesis from ground-level to aerial imagery, significant challenges remain.
+The drastic viewpoint differences between street and satellite views lead to distinct object appearances, resulting in substantial domain gaps across image types.
+As shown in Figure 1 (a), in cross-view scenarios, street-view images primarily capture ground and building facade details, while satellite views present more rooftop and macroscopic summary information.
+Even with state-of-the-art diffusion methods, the domain gap between views often results in generated images that are realistic yet not true to reality; the synthesized satellite imagery may differ significantly from the actual content of corresponding street-views, as shown in Figure 1(c).
+Another challenge is object occlusion in dense urban settings, where tall buildings or trees limit the visibility range in street-views, which is substantially smaller than the coverage of satellite images, as shown in Figure 1(b).
+Aerial images synthesized from a single street-view image struggle to maintain consistency with real satellite images in unseen areas.
+A effective approach is to leverage multiple neighboring street-view images for satellite view synthesis.
+However, this poses a challenge for the model in effectively utilizing geographic location information from street-views and integrating features from multiple street-view images.
+To address these challenges, this paper introduces SkyDiffusion, a ground-to-aerial image synthesis method based on diffusion models and the BEV paradigm, emulating a sky-view perspective.
+We first propose a Curved-BEV method, which transforms street-view images into Bird’s-Eye view to achieve domain alignment across viewpoints.
+Additionally, the Curved-BEV method introduces a multi-to-one mapping strategy, effectively integrating street-view images from different capture locations into a unified BEV space, thus overcoming the limited observation range in urban street-view images.
+Subsequently, we designed a BEV-guided diffusion model to generate realistic satellite images consistent with street-view content.
+Therefore, SkyDiffusion effectively overcomes the challenges of cross-view domain gaps and the limited observation range in densely occluded urban scenes, generating more realistic satellite images.
+Moreover, existing cross-view datasets, such as CVUSA [40] and VIGOR[46], were originally designed for cross-view retrieval tasks and thus offer limited diversity for cross-view synthesis tasks.
+We introduce an additional dataset, Ground2Aerial-3 (G2A-3), which includes three novel ground-to-aerial scene synthesis tasks: aerial image synthesis for disaster scenes based on Florida hurricane data, high-resolution historical aerial image generation (2007-2014), and low-altitude UAV image generation based on the MatrixCity [14] virtual urban environment.
+Ground2Aerial-3 provides practical value and significance for exploring cross-view synthesis tasks across more diverse scenarios.
+Our main contributions are as follows:
+  • We introduce SkyDiffusion, a novel ground-to-aerial synthesis method leveraging diffusion models and BEV paradigm to generate realistic, consistent aerial images.
   • We design a Curved-BEV method to transform street-view images into satellite views for domain alignment.
 It also includes "Multi-to-One" mapping strategy to enhance BEV perception range in densely occluded urban areas.
-  • SkyDiffusion outperforms state-of-the-art methods on suburban (CVUSA, CVACT) and urban (VIGOR-Chicago) cross-view datasets, with a 13.96% SSIM increase and 20.54% FID reduction.
-Related Work
-Cross-view image synthesis involves generating realistic scenes images with significant viewpoint changes, primarily divided into satellite-to-street (Shi et al.
-2022; Lu et al.
-2020; Qian et al.
-2023) and street-to-satellite generation (Regmi and Borji 2018, 2019).
-In this field, synthesizing satellite images from street views is a challenging yet important task.
-Regmi and Borji (2018) attempted to synthesize aerial views from a single street image, but the significant domain differences led to inconsistencies.
+  • We introduce Ground2Aerial-3, a new ground-to-aerial image synthesis dataset, featuring disaster scene aerial image, historical high-resolution satellite image, and low-altitude UAV image synthesis tasks.
+  • SkyDiffusion outperforms state-of-the-art methods on cross-view datasets across natural (CVUSA), suburban (CVACT), urban (VIGOR-Chicago), and various application scenarios (G2A-3), with an SSIM increase of 8.67% and a FID reduction of 21.50%.
+2 Related Work
+Cross-view image synthesis involves generating realistic scenes images with significant viewpoint changes, primarily divided into aerial-to-ground [29, 17, 22] and ground-to-aerial generation [23, 24].
+In this field, synthesizing satellite images from street-views is a challenging yet important task.
+Regmi and Borji [23] attempted to synthesize aerial views from a single street image, but the significant domain differences led to inconsistencies.
 Tang et al.
-(2019) used street-view images and aerial semantic maps for image-to-image translation to synthesize target satellite images, but semantic maps are often unavailable in real scenarios.
-Additionally, the performance of GANs has also limited the quality of street-to-satellite image synthesis in previous studies.
-Figure 2: Overview of the proposed SkyDiffusion framework, including the curved BEV transformation and BEV-controlled diffusion model with light manipulation.
-BEV transformation is a perspective domain conversion technique similar to polar transformation (Shi et al.
-2019; Toker et al.
-2021) and geo-transformation (Lu et al.
-2020), widely used in retrieval (Ye et al.
-2024b), navigation (Li et al.
-2022) and localization tasks (Song et al.
-2023; Wang et al.
-2023; Shi et al.
-2023b).
-Traditional BEV transformation techniques, which rely on camera parameters (Sarlin et al.
-2023) or depth estimation methods (Teng et al.
-2024; Ye et al.
-2024a), face challenges such as the difficulty in obtaining camera parameters and the high computational cost of depth estimation.
-Recently, some methods have been proposed to map street-view images to the BEV perspective using geometric relations at the image (Wang et al.
-2023) or feature level (Shi et al.
-2023b).
+[33] used street-view images and aerial semantic maps for image-to-image translation to synthesize target satellite images, but semantic maps are often unavailable in real scenarios.
+Additionally, the performance of GANs has also limited the quality of ground-to-aerial image synthesis in previous studies.
+Concurrent work [1] also synthesizes satellite images using text guidance and staged BEV semantic segmentation results, but these methods require additional text data and an extra semantic segmentation phase on street-view images.
+Figure 2: Overview of the proposed SkyDiffusion framework.
+It include the curved BEV transformation and BEV-controlled diffusion model with light manipulation.
+The lower parts present the results of One-to-One and Multi-to-One BEV transformations, respectively.
+BEV transformation is a perspective domain conversion technique similar to polar transformation [28, 35] and geo-transformation [17], widely used in retrieval [42], navigation [15] and localization tasks [32, 39, 30].
+Traditional BEV transformation techniques, which rely on camera parameters [27] or depth estimation methods [34, 41], face challenges such as the difficulty in obtaining camera parameters and the high computational cost of depth estimation.
+Recently, some methods have been proposed to map street-view images to the BEV perspective using geometric relations at the image [39] or feature level [30].
 However, such methods are based on the ground plane assumption, where all mapped points are located below the camera, limiting the mapping of upper street scene features and causing information loss.
-Furthermore, to our knowledge, no studies have yet explored the use of BEV methods for street-to-satellite image synthesis task.
+Moreover, most existing panoramic BEV transformation methods primarily focus on one-to-one mappings, with relatively little research on multi-to-one BEV joint mappings.
 Novel view synthesis based on diffusion models.
-Diffusion models have shown significant potential in recent studies on novel view synthesis (Kawar et al.
-2023; Gao et al.
-2024; Tseng et al.
-2023).
-DreamBooth (Ruiz et al.
-2023) and AerialDiffusion (Kothandaraman et al.
-2023a) utilize text guidance to achieve top-down image generation by fine-tuning diffusion models.
+Diffusion models have shown significant potential in recent studies on novel view synthesis [7, 4, 37].
+DreamBooth [26] and AerialDiffusion [9] utilize text guidance to achieve top-down image generation by fine-tuning diffusion models.
 However, it is challenging to fully encapsulate the complexity of real-world scenes using textual information alone.
-Zero-1-to-3 (Liu et al.
-2023) and Zero 123++ (Shi et al.
-2023a) use multi-view images and camera poses to guide diffusion models in learning object structures and textures, enabling new viewpoint synthesis.
-However, these methods are limited to object-level tasks with strong viewpoint correlations and struggle to handle the significant viewpoint changes in cross-view generation.
-Besides, diffusion-based image-to-image translation methods can also generate realistic satellite images (Li et al.
-2023a; Brooks, Holynski, and Efros 2023; Zhang, Rao, and Agrawala 2023).
+Besides, diffusion-based image-to-image translation methods can also generate realistic satellite images [12, 2, 43].
 However, the substantial domain gap between street-view images and satellite images presents a significant challenge in generating content-consistent satellite images.
-Methods
+3 Methods
 As illustrated in Fig.
 2, this paper introduces SkyDiffusion, a novel method for synthesizing satellite images from corresponding street-view images.
-It initially applies a curved BEV transformation to the input street-view images, converting the perspective to a top-down view using either one-to-one or multi-to-one mappings.
-Subsequently, the BEV-Controlled diffusion model with light manipulation is employed for the controlled synthesis of satellite images.
-Following this, we detail our network training process.
-Curved BEV Transformation
+It initially applies a curved BEV transformation to the input street-view images, converting the perspective to a top-down view using either one-to-one or multi-to-one mappings (Section 3.1).
+Subsequently, a BEV-guided diffusion model controls the synthesis of satellite images, with an optional light manipulation module included(Section 3.2).
+Following this, we detail our network training process (Section 3.3).
+3.1 Curved BEV Transformation
 One-to-One BEV.
-Existing BEV methods based on ground plane assumptions ($P(x,y,z=0)$) and geometric relationships achieve the mapping from street view to the BEV plane(Shi et al.
-2023b; Wang et al.
-2023).
+Existing BEV methods based on ground plane assumptions ($P(x,y,z=0)$) and geometric relationships achieve the mapping from street-view to the BEV plane[30, 39].
 However, this assumption fails to map the content above the street-view image and results in noticeable distortion as points move away from the center area.
 Thus, we propose an enhanced curved BEV transformation to better capture cross-view information and address this issue.
-Figure 3: Schematic of the curved BEV transformation.
-The upper part (a) details the transformation process, highlighting the two mapping points at the top and bottom of the street view.
-The lower parts (b) and (c) present the results of One-to-One and Multi-to-One BEV transformations, respectively.
 As shown in Fig.
-3 (a), we use $P(i,j)$ to denote an index on the BEV space, and $P(x,y,z)$ to illustrate a point in a three-dimensional coordinate system with the camera on the z-axis.
+3, we use $P(i,j)$ to denote an index on the BEV space, and $P(x,y,z)$ to illustrate a point in a three-dimensional coordinate system with the camera on the z-axis.
 We assume that the BEV plane is an upward curved surface, with the height (z-axis) increasing rapidly as the 3D points move away from the center (the central part of Fig 3(a)).
@@
 Furthermore, we can convert $P(x,y,z)$ to spherical coordinates $P(\theta,\varphi)$.
-Based on the equirectangular projection properties of street view images, $P(\theta,\varphi)$ corresponds to the panorama image index $P(u,v)$.
+Based on the equirectangular projection properties of street-view images, $P(\theta,\varphi)$ corresponds to the panorama image index $P(u,v)$.
 This allows us to establish the mapping relationship between $P(x,y,z)$ and $P(u,v)$ from the following equation, where $H$ represents the camera height, and $w$, $h$ represent the width and height of the street image.
@@
 1 and Eq.
-2, we achieve the projection from the street-view image to the bird’s eye view without requiring camera calibration parameters or depth estimation methods (Teng et al.
-2024; Ye et al.
-2024a) (Fig.
-3 (b)).
-This effectively maps upper street view features like distant roads and buildings.
+2, we achieve the projection from the street-view image to the bird’s eye view without requiring camera calibration parameters or depth estimation methods [34, 41].
+This effectively maps upper street-view features like distant roads and buildings.
 Since the mapping relationship for each transformation is fixed, the computational overhead is minimal.
+Figure 3: Schematic of the curved BEV transformation. It illustrates the mapping of two points on the BEV plane to the top and bottom of the street-view image during transformation.
 Multi-to-One BEV.
-Due to the limitations imposed by dense buildings on street view observations, the BEV sensing range is very limited in urban datasets like VIGOR.
-To address this, we propose a "Multi-to-one" BEV mapping method that unifies multiple street view BEV mappings results $\textit{BEV}_{\mathrm{cam}_{k}}$ into a single satellite coordinate system based on the relationships of camera position (Fig.
-3 (c)).
-We use a simple semantic segmentation network to identify and discard the regions with severe BEV distortion, typically those obstructed by buildings.
-We select the street view mapping result closest to the capture point for overlapping areas following Eq.
+Due to the limitations imposed by dense buildings on street-view observations, the BEV sensing range is very limited in urban datasets like VIGOR.
+To address this, we propose a "Multi-to-one" BEV mapping method that unifies multiple street-view BEV mappings results $\textit{BEV}_{\mathrm{cam}_{k}}$ into a single satellite coordinate system based on the relationships of camera position.
+We select the street-view mapping result closest to the capture point for overlapping areas following Eq.
 3, where $\textit{BEV}\left(x,y\right)$ represents the unified BEV space corresponding to the satellite image, $x,y$ corresponds to the index position of the BEV plane, $\Delta x_{k}$ and $\Delta y_{k}$ represent the offset of the street-view image from the satellite center, and $x_{\mathrm{cam}_{i}}$ and $y_{\mathrm{cam}_{i}}$ represent the coordinates of camera position in the satellite coordinate system.
@@
 This enables us to synthesize large-scale scenes using street-view images with limited local information.
-BEV-Controlled Synthesis with Light Manipulation
+3.2 BEV-Controlled Diffusion Model
 Curved BEV controlled synthesis.
@@
 The denoising process can occur in pixel space or in a latent space encoded from training data.
-Inspired by ControlNet (Zhang, Rao, and Agrawala 2023), we enhance the control over the generated aerial views by replicating the SD module, enabling the transformed BEV images to effectively guide the synthesis process.
+Inspired by ControlNet [43], we enhance the control over the generated aerial views by replicating the SD module, enabling the transformed BEV images to effectively guide the synthesis process.
 Specifically, we first use a small encoder to capture the cross-view associative information of the input BEV image $I_{bev}$ such as road directions and building positions, to achieve BEV image feature embedding $c_{bev}$.
@@
 Guided by the BEV image, the diffusion model incorporates cross-view associative information, synthesizing aerial views with consistent content.
-Light Manipulation.
+Optional Light Manipulation.
 Given the differences in perspectives and timing between satellite and street images, there is no direct correlation in lighting conditions.
 To address this, we designed a light manipulation module to make the lighting conditions of the synthesized satellite images more adaptable.
-We employ the DNCM for modeling deterministic color mapping (Ke et al.
-2023).
+We employ the DNCM for modeling deterministic color mapping [8].
 Initially, we use nDNCM to obtain the "image content information" $I_{c}$ from the preliminary synthesis results $I_{0}$ of the satellite images.
@@
 {$&I_{s}=s\textit{DNCM}(\textit{nDNCM}(I_{0},I_{c}),r_{s})\\ &\textit{DNCM}(I,T)=(I_{h\times w,3}\cdot P_{3,k})\cdot T_{k,k}\cdot Q_{k,3}\end{aligned}\right.$ (5)
-Figure 4: Qualitative Comparison of Synthesis Results.
-The first and second rows present the results on the CVUSA dataset, while the third and fourth rows display the results on the CVACT dataset.
-The fifth and sixth rows showcase the visualized outcomes on the VIGOR-Chicago dataset.
 $I_{c}$ and $r_{s}$ are the content information and color reference information extracted by the encoder $E$, as shown in Fig.
@@
 We obtain lighting information from randomly selected satellite images in the training set, and standard images can also be used as lighting control conditions.
-Network Training
+3.3 Network Training
 Loss Function.
@@
 It then learns a network $\epsilon_{\theta}$ to predict the noise added to the noisy satellite image $Z_{t}$ based on the given BEV image as the conditional input.
-Since text cannot fully describe the street view scene, our cross-view image synthesis is essentially an image-conditioned synthesis task, with text not being our optimization target.
+Since text cannot fully describe the street-view scene, our cross-view image synthesis is essentially an image-conditioned synthesis task, with text not being our optimization target.
 This helps the model focus on generating satellite images based on the BEV conditions.
@@
 The loss $\mathcal{L}$ is the learning objective of the diffusion model, where $t$ represents the time step.
-Following previous work (Ke et al.
-2022, 2023), we apply different LUT filters to the satellite images to simulate the results under various lighting conditions for training the light manipulation module.
-During training, we use L2 consistency loss $\mathcal{L}_{\text{con}}$ to maintain image content consistency and L1 reconstruction loss $\mathcal{L}_{\text{rec}}$ to learn color style transfer.
+Implementation Details.
+We implement SkyDiffusion using the pre-trained Stable Diffusion v1.5 model [25], with an unlocked diffusion decoder and a classifier-free guidance scale of 9.0 [6].
+Final inference sampling uses 50 steps with the DDIM strategy [31].
+Training is on eight NVIDIA A100 GPUs with a batch size of 128 for 100 epochs.
+For optional light manipulation, the reference satellite image is randomly selected from normally toned images in the training set, and the DNCM mapping matrix parameter $k$ is set to 32 (details in Supplementary Material).
+4 Dataset
+We propose Ground2Aerial-3, a multi-task cross-view synthesis dataset designed to explore the performance of cross-view synthesis methods in several novel scenarios.
+As shown in Figure 3, G2A-3 contains nearly 10k pairs of street-view and aerial images, covering disaster scene aerial image synthesis, historical high-resolution satellite image synthesis, and low-altitude UAV image synthesis.
+The dataset of each task is randomly split into training and test sets with a ratio of 5:1.
+The ground street-view images are 1024×512, with true north aligned at the center, and the aerial images are 512×512, aligned with the center of the ground images.
+4.1 Disaster Scene Aerial image
+Both satellite and aerial images are limited by their acquisition frequency, which imposes certain constraints on rapid response to natural disasters.
+Synthesizing aerial images based on street-view images captured by ground cameras can facilitate fast disaster assessment and analysis.
+CVIAN [13] is a cross-view dataset containing street-view and high-resolution satellite imagery in Florida, USA, following Hurricane Ian in 2022.
+From CVIAN’s street-view and satellite data, we filtered 2.7k image pairs of moderately or severely affected areas and adjusted the north direction of street-view images to the central column, preparing the data for the disaster scene aerial image generation task.
+Figure 4: Illustration of the Ground2Aerial-3 dataset.
+4.2 Historical High-Resolution Satellite image
+Obtaining high-resolution satellite imagery, particularly ultra-high resolution images exceeding 0.3m, has been challenging.
+It wasn’t until the recent launch of WorldView-3 (2014) and WorldView-4 (2018) that high-resolution coverage at near-global scales became available.
+Previously, ultra-high resolution images were often limited to small areas captured by aerial photography, with both spatial and temporal coverage constraints.
+Synthesizing satellite images based on historical street-view images can help fill this gap, providing additional complementary information.
+We collected historical street-view images from 2007 to 2014 on the Mapillary  [^1] https://www.mapillary.com/ platform and used corresponding aerial imagery from the massGIS platform  [^2] https://www.mass.gov/info-details/massgis-data-layers as ground truth to conduct high-resolution satellite image synthesis tasks.
+4.3 Low-Altitude UAV Image
+Low-altitude UAV imagery, with ultra-high resolution below 0.05 meters, has potential applications in tasks such as lane data collection for autonomous driving.
+However, the coverage and cost of UAV imagery are limited, and safety regulations further restrict coverage range and update frequency.
+Synthesizing low-altitude UAV images from more accessible street-view images presents a potential application scenario.
+Using virtual MatrixCity [14] data, we employed the UE engine to set up six single-view cameras at ground level, rendering and stitching them into panoramic street-view images.
+At the same locations, we positioned downward-facing cameras at an altitude of 20 meters to render low-altitude UAV imagery, creating a cross-view dataset for low-altitude UAV image synthesis tasks.
+Figure 5: Qualitative comparison of different methods synthesis results on three datasets.
+5 Experiments
+In this section, we introduce the datasets utilized and the experimental setup.
+We then conduct both qualitative and quantitative comparisons of SkyDiffusion with state-of-the-art cross-view synthesis methods.
+Finally, we perform ablation studies to evaluate the effectiveness of each module.
+5.1 Experimental Setup
+Dataset Details.
+CVUSA [40] and CVACT [16] are two commonly used cross-view datasets, containing one-to-one, center-aligned aerial and ground street-view images.
+They mainly covers suburban scenes, with low-rise buildings and open views.
+The CVACT dataset is regionally segmented.
+Although CVUSA uses random partitioning, its data distribution is sparse, with no overlap among satellite images.
+VIGOR-Chicago [46] is a subset of the VIGOR cross-view dataset, containing the Chicago area.
+In this dataset, street-view panoramas and satellite images are not center-aligned, with multiple street-view images covering the same satellite image region, forming a multi-to-one mapping relationship.
+We modified the train-test split to a 4:1 ratio and ensured that the same satellite images do not appear in both the training and testing sets to prevent information leakage.
+The G2A-3 dataset has been described in Section 4.
+All input street images are set to 512 $\times$ 1024 pixels, with the center column aligned to the north.
+Unlike previous methods that synthesize low-resolution satellite images (64 $\times$ 64), this experiment aims to meet practical needs by generating higher-resolution satellite images (512 $\times$ 512 pixels).
+In the Curved-BEV transformation, $\lambda$ is set to 2 and $H$ to 1.5 for CVUSA, CVACT and G2A-3 datasets.
+For the VIGOR-Chicago dataset, when performing a one-to-one mapping, $\lambda$ is set to 10; for a Multi-to-One mapping, $\lambda$ is set to 0, the primary focus is on retaining the street-view central location information.
+The Curved-BEV transformation generates results that are size-matched with satellite images.
+Comparison Methods and Metric.
+We compared our method with several state-of-the-art cross-view synthesis methods, including the commonly used cross-view satellite synthesis methods X-Seq [23], SelectionGAN (without semantic masks) [33], and CDTE [35], image translation methods such as the GAN-based CUT [20] and diffusion-based Instruct pix2pix (Instr-p2p) [2], ControlNet [43] and img2img-turbo [21], as well as novel view synthesis methods like Aerial Diff [9] and HawkI [10].
+Except for the fine-tuned diffusion model, which requires brief textual descriptions, all other models use the same street-view input.
+For a fair comparison, the results in the table 1 are presented without the use of the Optional Light Manipulation module for Skydiffusion.
+Following previous work [17, 35, 23], we employ commonly used metrics such as SSIM, PSNR and LPIPS to evaluate the content consistency of the synthesized images, and use FID [5] to assess the realism of images.
 Method CVUSA CVACT VIGOR-Chicago
 FID ($\downarrow$) SSIM ($\uparrow$) PSNR ($\uparrow$) LPIPS ($\downarrow$) FID ($\downarrow$) SSIM ($\uparrow$) PSNR ($\uparrow$) LPIPS ($\downarrow$) FID ($\downarrow$) SSIM ($\uparrow$) PSNR ($\uparrow$) LPIPS ($\downarrow$)
-X-Seq 161.16 0.084 11.97 0.706 190.12 0.042 12.41 0.661 - - - -
-CDTE 122.84 0.143 10.01 0.694 160.81 0.091 12.59 0.663 - - - -
-SelectionGAN 116.57 0.129 12.38 0.742 100.21 0.116 12.04 0.684 149.53 0.127 11.77 0.778
-Aerial Diff 136.18 0.103 10.06 0.855 127.29 0.108 10.24 0.878 123.16 0.141 11.49 0.831
-HawkI 90.98 0.114 11.09 0.783 135.75 0.112 10.97 0.764 123.47 0.124 9.67 0.752
-CUT 72.83 0.121 12.09 0.687 62.22 0.102 13.11 0.664 69.42 0.169 13.66 0.665
-img2img-turbo 77.95 0.127 12.64 0.685 73.24 0.103 12.19 0.679 80.10 0.135 11.28 0.690
-Instr-p2p 38.01 0.138 12.29 0.697 49.62 0.116 12.51 0.682 50.30 0.163 10.62 0.689
-Ours 29.18 0.168 14.58 0.635 35.51 0.123 13.58 0.657 45.29 0.186 11.69 0.661
+X-Seq [23] 161.16 0.084 11.97 0.706 190.12 0.042 12.41 0.661 - - - -
+CDTE [35] 122.84 0.143 10.01 0.694 160.81 0.091 12.59 0.663 - - - -
+SelGAN [33] 116.57 0.129 12.38 0.742 100.21 0.116 12.04 0.684 149.53 0.127 11.77 0.778
+Aerial Diff [9] 136.18 0.103 10.06 0.855 127.29 0.108 10.24 0.878 123.16 0.141 11.49 0.831
+HawkI [10] 90.98 0.114 11.09 0.783 135.75 0.112 10.97 0.764 123.47 0.124 9.67 0.752
+CUT [20] 72.83 0.121 12.09 0.687 62.22 0.102 13.11 0.664 69.42 0.169 13.66 0.665
+I2I-Turbo [21] 77.95 0.127 12.64 0.685 73.24 0.103 12.19 0.679 80.10 0.135 11.28 0.690
+Instr-p2p [2] 38.01 0.138 12.29 0.697 49.62 0.116 12.51 0.682 50.30 0.163 10.62 0.689
+ControlNet [43] 32.45 0.149 12.63 0.650 62.21 0.115 11.95 0.682 53.27 0.170 10.38 0.666
+Ours 29.18 0.168 14.58 0.635 36.48 0.118 12.85 0.645 45.29 0.186 11.69 0.661
 Table 1: Quantitative comparison of different methods on CVUSA, CVACT and VIGOR-Chicago.
-Implementation Details.
-We implement SkyDiffusion using the pre-trained Stable Diffusion v1.5 model (Rombach et al.
-2022), with an unlocked diffusion decoder and a classifier-free guidance scale of 9.0 (Ho and Salimans 2022).
-Final inference sampling uses 50 steps with the DDIM strategy (Song, Meng, and Ermon 2021).
-Training is on eight NVIDIA A100 GPUs with a batch size of 128 for 100 epochs.
-For light manipulation, the reference satellite image is randomly selected from normally toned images in the training set, and the DNCM mapping matrix parameter $k$ is set to 32 (details in Supplementary Material).
-Experiments
-In this section, we introduce the three datasets utilized and the experimental setup.
-We then conduct both qualitative and quantitative comparisons of SkyDiffusion with state-of-the-art cross-view synthesis methods.
-Finally, we perform ablation studies to evaluate the effectiveness of each module.
-Datasets
-CVUSA & CVACT (Suburban dataset).
-CVUSA (Workman, Souvenir, and Jacobs 2015) and CVACT (Liu and Li 2019) are two widely used cross-view datasets.
-They contain centrally aligned satellite and street-view images from regions in the United States and Canberra, Australia, respectively, with a one-to-one mapping relationship.
-The data primarily consists of suburban scenes characterized by relatively low-rise buildings and open views.
-The CVUSA dataset’s training and testing sets are randomly split, while the CVACT dataset is split by region.
-VIGOR-Chicago (Urban dataset).
-VIGOR-Chicago (Zhu, Yang, and Chen 2021) is a subset of the VIGOR dataset, containing cross-view image pairs collected from Chicago urban area.
-In this dataset, the street panorama and satellite images are not centrally aligned, and multiple street-view images cover the same satellite image area, with overlapping regions between different satellite images, forming a multi-to-one mapping relationship.
-All input street images are set to 512 × 1024 pixels, with the center column aligned to the north.
-Unlike previous methods that synthesize low-resolution satellite images (64 $\times$ 64 or 256 $\times$ 256 pixels), this experiment aims to meet practical needs by generating higher-resolution satellite images (512 $\times$ 512 pixels).
-In the Curved-BEV transformation, $\lambda$ is set to 3 and $H$ to 1.5 for CVUSA and CVACT datasets.
-For the VIGOR-Chicago dataset, $\lambda$ is set to 10, with $H$ unchanged.
-The Curved-BEV transformation results in street images that match the size of the satellite images.
-Experimental Setup
-We compared our method on three datasets with several state-of-the-art cross-view synthesis methods, including the commonly used cross-view satellite synthesis methods X-Seq (Regmi and Borji 2018), SelectionGAN (without semantic masks) (Tang et al.
-2019), and CDTE (Toker et al.
-2021), image translation methods such as the GAN-based CUT (Park et al.
-2020) and diffusion-based Instruct pix2pix (Instr-p2p) (Brooks, Holynski, and Efros 2023) and img2img-turbo (Parmar et al.
-2024), as well as novel view synthesis methods like Aerial Diff (Kothandaraman et al.
-2023a) and HawkI (Kothandaraman et al.
-2023b).
-Except for the fine-tuned diffusion model, which requires brief textual descriptions, all other models use the same street-view input for a fair comparison.
-Following previous work (Lu et al.
-2020; Toker et al.
-2021; Regmi and Borji 2018), we employ commonly used metrics such as SSIM, PSNR and LPIPS to evaluate the content consistency of the synthesized images, and use FID (Heusel et al.
-2017) to assess the realism of images.
-Comparison with State-of-the-Art Methods
-On the suburban CVUSA and CVACT datasets, our SkyDiffusion method achieved the outstanding results, as shown in Table 1.
-Compared to state-of-the-art methods, it reduced FID by 25.83% and increased SSIM by 13.89%, demonstrating its superiority in synthesizing realistic and consistent satellite images.
+Method FID ($\downarrow$) SSIM ($\uparrow$) PSNR ($\uparrow$) LPIPS ($\downarrow$)
+Disaster Scene image synthesis
+ControlNet 138.11 0.183 12.95 0.690
+Ours 114.64 0.188 13.02 0.694
+Historical aerial image synthesis
+ControlNet 36.48 0.117 12.85 0.665
+Ours 35.51 0.123 13.58 0.657
+Low-altitude UAV image synthesis
+ControlNet 52.76 0.371 16.55 0.598
+Ours 32.11 0.452 16.86 0.390
+Table 2: Quantitative comparison of methods on G2A-3 Dataset.
+5.2 Comparison with State-of-the-Art Methods
+On the suburban CVUSA and CVACT datasets, SkyDiffusion achieved the outstanding results, as shown in Table 1.
+Compared to state-of-the-art methods, it reduced FID by 25.72% and increased SSIM by 7.68%, demonstrating its superiority in synthesizing realistic and consistent satellite images.
 The visual results in Fig.
@@
 Additional visual results can be found in the supplementary materials.
-In the urban VIGOR-Chicago dataset, SkyDiffusion reduced FID by 9.96% and improved SSIM by 14.11% compared to the state-of-the-art method, as shown in Table 1.
-For fair comparison, the VIGOR results for SkyDiffusion use only a single street-view image.
+Figure 6: Qualitative Visualization Results on the G2A-3 Dataset.
+In the urban VIGOR-Chicago dataset, our method reduced FID by 14.98% and improved SSIM by 9.41% compared to the state-of-the-art method, as shown in Table 1.
+The VIGOR results for SkyDiffusion use only a single street-view image.
 The results using Multi-to-One BEV generation with multiple street-view images can be found in the ablation study.
@@
 While the non-associative information on building rooftops cannot be consistently generated, it still produces reasonable and realistic results.
-Ablation Study
+Figure 7: Qualitative comparison of synthesis results for One-to-One and Multi-to-one BEV.
+As shown in Table 2, the tasks on the G2A-3 dataset present certain challenges; however, our method achieves significant performance improvements over the commonly used image-conditioned synthesis method, ControlNet.
+SkyDiffusion reduces the FID by an average of 19.60% and increases the SSIM by an average of 9.90%.
+The visual results in Figure 6 demonstrate SkyDiffusion’s ability to synthesize various types of aerial images across different G2A-3 task settings.
+In Figure 6(a), prominent disaster areas in the street-view image (orange box indicating debris of damaged buildings and red box showing trees downed by the hurricane) are clearly visible in the synthesized satellite image, aiding in subsequent disaster area assessment and analysis.
+Figure 6(b) shows that SkyDiffusion can synthesize realistic early historical imagery, effectively filling gaps where early high-resolution aerial images are lacking.
+Figure 6(c) highlights SkyDiffusion’s exceptional performance in low-altitude UAV image synthesis.
+Due to the lower altitude of UAV captures, there is greater overlap and closer domain similarity between scenes, resulting in superior synthesis quality.
+SkyDiffusion demonstrates excellent aerial image synthesis capabilities, advancing potential applications in various aerial image synthesis scenarios.
+5.3 Ablation Study and More Exploration
 Curved-BEV controlled synthesis.
@@
 Additional ablation study results can be found in the supplementary materials.
-Method FID ($\downarrow$) SSIM ($\uparrow$) PSNR ($\uparrow$) LPIPS ($\downarrow$)
-CVUSA
-Baseline 32.45 0.149 12.63 0.650
-Ours (C-BEV) 29.18 0.168 14.58 0.635
-CVACT
-Baseline 62.21 0.115 11.95 0.682
-Ours (C-BEV) 35.51 0.123 13.58 0.657
-VIGOR-Chicago
-Baseline 53.27 0.170 10.38 0.666
-Ours (C-BEV) 45.29 0.186 11.69 0.661
-Ours (C-BEV Multi) 31.90 0.205 11.15 0.651
-Table 2: Ablation study of the Curved-BEV module. "Baseline" represents directly using street-view image,"C-BEV" denotes using Curved-BEV transformation, and "Multi" stands for Multi-to-One strategy.
-Figure 5: Qualitative comparison of synthesis results for One-to-one and Multi-to-one BEV.
+C-BEV Light FID ($\downarrow$) SSIM ($\uparrow$) PSNR ($\uparrow$) LPIPS ($\downarrow$)
+CVACT 62.21 0.115 11.95 0.682
+36.48 0.118 12.85 0.645
+35.51 0.123 13.58 0.657
+VIGOR 53.27 0.170 10.38 0.666
+45.29 0.186 11.69 0.661
+(Multi) 37.90 0.193 10.39 0.660
+(Multi) 31.90 0.205 11.15 0.651
+Table 3: Ablation study of the Curved-BEV and optional Light Manipulation Modules. "Multi" stands for Multi-to-One strategy, and "Light" represents using light manipulation.
 Light manipulation.
-The ablation experiments in Table 3 indicate that the Light Manipulation module aligns the lighting conditions of the synthesized images with those of the target domain images, improving SSIM and PSNR metrics.
-This is because satellites typically capture large areas with relatively uniform overall brightness.
-Using a small satellite image slice as a reference effectively captures the lighting conditions at the time of capture.
-Due to the dispersed distribution of the CVUSA data and the more complex lighting conditions, the improvements are not as pronounced as on other datasets.
-The light processing module maintains the original image content while providing more flexible lighting conditions for the synthesized satellite images.
-More visualization results can be found in the supplementary material.
-Method FID ($\downarrow$) SSIM ($\uparrow$) PSNR ($\uparrow$) LPIPS ($\downarrow$)
-CVUSA
-Ours (w/o light) 29.02 0.167 14.07 0.643
-Ours 29.18 0.168 14.58 0.635
-CVACT
-Ours (w/o light) 36.48 0.117 12.85 0.665
-Ours 35.51 0.123 13.58 0.657
-VIGOR-Chicago
-Ours (w/o light) 46.77 0.179 10.46 0.666
-Ours 45.29 0.186 11.69 0.661
-Table 3: Ablation study of the light manipulation module.
-Figure 6: Synthesis results for cross-dataset generalization.
-Method FID ($\downarrow$) SSIM ($\uparrow$) PSNR ($\uparrow$) LPIPS ($\downarrow$)
-CVACT $\rightarrow$ VIGOR-Chicago
-Instr-p2p 100.80 0.152 9.14 0.717
-Ours (w/o light) 89.76 0.155 9.45 0.739
-Ours 63.04 0.162 11.24 0.723
-VIGOR-Chicago $\rightarrow$ CVACT
-Instr-p2p 121.29 0.113 8.90 0.759
-Ours (w/o light) 107.58 0.116 8.76 0.751
-Ours 100.58 0.121 11.36 0.714
-Table 4: Cross-dataset generalization assessment.
-Cross-Dataset Generalization
-We trained the model on CVACT and tested it on the VIGOR-Chicago, and vice versa, to evaluate cross-dataset generation capability, as shown in Table 4.
-Compared to Instruct pix2pix, our method (w/o light) demonstrates superior performance across metrics.
-Figure 6 shows the synthesis results for cross-dataset generalization.
-Our method effectively preserves scene content such as road directions and intersections.
-However, due to the distinct city styles of the CVACT and VIGOR datasets and the fact that the model was trained on a single dataset, there are noticeable differences in elements such as buildings and lighting conditions.
-The aim of this task is not to generate identical satellite images but to create scenes with consistent content layouts while exhibiting distinctly different styles.
-Additionally, we can flexibly use the light manipulation module to obtain light information from a random image in the target dataset’s training set, better matching the target domain’s lighting conditions.
-Conclusion
+Aerial and satellite images typically capture large areas with relatively uniform overall brightness.
+Using small aerial image patches as references effectively captures the lighting conditions at the time of capture.
+We applied an optional Light Manipulation module to align the lighting conditions of the synthesized image with those of the target domain image, as shown in Table 3 The results indicate that Light Manipulation can effectively improve SSIM and PSNR metrics.
+This module preserves the original image content while providing more flexible lighting conditions for the synthesized satellite image.
+Additional visualizations can be found in the supplementary material.
+6 Conclusion
 In this study, we introduce SkyDiffusion, a novel approach specifically designed for street images to satellite images cross-view synthesis.
 SkyDiffusion operates solely with street images as input, utilizing a BEV Paradigm and diffusion models to generate satellite images.
-SkyDiffusion achieves state-of-the-art performance in both content consistency and image realism on suburban and urban cross-view datasets, demonstrating its superior capabilities in cross-view synthesis tasks.
-The satellite and street-view domains bridging method, Curved-BEV, within this approach also holds valuable insights for tasks such as cross-view retrieval and geolocation.
-In the future work, we plan to extend our method to more cities and further refine it for specific downstream applications, such as urban planning and virtual tourism.
+SkyDiffusion achieves state-of-the-art performance in both content consistency and image realism on across multiple cross-view datasets, demonstrating its superior capabilities.
+Additionally, we introduce a cross-view synthesis dataset, Ground2Aerial-3, featuring aerial image synthesis tasks for multiple new scenes, providing practical value and inspiration for future cross-view synthesis research.

### v4

@@
-SkyDiffusion: Ground-to-Aerial Image Synthesis with Diffusion Models and BEV Paradigm
+1Association for the Advancement of Artificial Intelligence
 Abstract
-Ground-to-aerial image synthesis focuses on generating realistic aerial images from corresponding ground street view images while maintaining consistent content layout, simulating a top-down view.
-The significant viewpoint difference leads to domain gaps between views, and dense urban scenes limit the visible range of street views, making this cross-view generation task particularly challenging.
-In this paper, we introduce SkyDiffusion, a novel cross-view generation method for synthesizing aerial images from street view images, utilizing a diffusion model and the Bird’s-Eye View (BEV) paradigm.
-The Curved-BEV method in SkyDiffusion converts street-view images into a BEV perspective, effectively bridging the domain gap, and employs a "multi-to-one" mapping strategy to address occlusion issues in dense urban scenes.
-Next, SkyDiffusion designed a BEV-guided diffusion model to generate content-consistent and realistic aerial images.
-Additionally, we introduce a novel dataset, Ground2Aerial-3, designed for diverse ground-to-aerial image synthesis applications, including disaster scene aerial synthesis, historical high-resolution satellite image synthesis, and low-altitude UAV image synthesis tasks.
-Experimental results demonstrate that SkyDiffusion outperforms state-of-the-art methods on cross-view datasets across natural (CVUSA), suburban (CVACT), urban (VIGOR-Chicago), and various application scenarios (G2A-3), achieving realistic and content-consistent aerial image generation.
-More result and dataset information can be found at https://opendatalab.github.io/skydiffusion/.
-1 Introduction
-In fields such as land cover classification [18, 36], urban planning [47, 45], and disaster response [11, 44, 38], the analysis of aerial imagery plays a crucial role.
+Street-to-satellite image synthesis focuses on generating realistic satellite images from corresponding ground street-view images while maintaining a consistent content layout, similar to looking down from the sky.
+The significant differences in perspectives create a substantial domain gap between the views, making this cross-view generation task particularly challenging.
+In this paper, we introduce SkyDiffusion, a novel cross-view generation method for synthesizing satellite images from street-view images, leveraging diffusion models and Bird’s Eye View (BEV) paradigm.
+First, we design a Curved-BEV method to transform street-view images to the satellite view, reformulating the challenging cross-domain image synthesis task into a conditional generation problem.
+Curved-BEV also includes a "Multi-to-One" mapping strategy for leveraging multiple street-view images within the same satellite coverage area, effectively solving the occlusion issues in dense urban scenes.
+Next, we design a BEV-controlled diffusion model to generate satellite images consistent with the street-view content, which also incorporates a light manipulation module to make the lighting conditions of the synthesized satellite images more flexible.
+Experimental results demonstrate that SkyDiffusion outperforms state-of-the-art methods on both suburban (CVUSA & CVACT) and urban (VIGOR-Chicago) cross-view datasets, with an average SSIM increase of 13.96% and a FID reduction of 20.54%, achieving realistic and content-consistent satellite image generation.
+The code and models will be made publicly available.
+Introduction
+In fields such as land cover classification [lv2023novel, tong2020land], urban planning [zhu2019understanding, zhao2019exploring], and disaster response [kucharczyk2021remote, zhao2024see, wang2023hierarchical], the analysis of aerial imagery plays a crucial role.
 The cross-view Ground-to-Aerial Image Synthesis task leverages ground street-view images to synthesize corresponding aerial images at the same location.
-Street-view images can be captured and uploaded by users on platforms such as Google Street View or Mapillary, offering high flexibility in data collection and frequent temporal coverage [1].
-Therefore, Ground-to-Aerial Image Synthesis can support rapid response during natural disasters [19] or fill historical gaps in high-resolution satellite imagery records [3], among other applications.
+Street-view images can be captured and uploaded by users on platforms such as Google Street View or Mapillary, offering high flexibility in data collection and frequent temporal coverage [arrabi2024cross].
+Therefore, Ground-to-Aerial Image Synthesis can support rapid response during natural disasters [ou2023method] or fill historical gaps in high-resolution satellite imagery records [cornebise2022open], among other applications.
 Figure 1: Illustration of the cross-view image synthesis task.
@@
 Aerial images synthesized from a single street-view image struggle to maintain consistency with real satellite images in unseen areas.
-A effective approach is to leverage multiple neighboring street-view images for satellite view synthesis.
+An effective approach is to leverage multiple neighboring street-view images for satellite view synthesis.
 However, this poses a challenge for the model in effectively utilizing geographic location information from street-views and integrating features from multiple street-view images.
@@
 Therefore, SkyDiffusion effectively overcomes the challenges of cross-view domain gaps and the limited observation range in densely occluded urban scenes, generating more realistic satellite images.
-Moreover, existing cross-view datasets, such as CVUSA [40] and VIGOR[46], were originally designed for cross-view retrieval tasks and thus offer limited diversity for cross-view synthesis tasks.
-We introduce an additional dataset, Ground2Aerial-3 (G2A-3), which includes three novel ground-to-aerial scene synthesis tasks: aerial image synthesis for disaster scenes based on Florida hurricane data, high-resolution historical aerial image generation (2007-2014), and low-altitude UAV image generation based on the MatrixCity [14] virtual urban environment.
-Ground2Aerial-3 provides practical value and significance for exploring cross-view synthesis tasks across more diverse scenarios.
+Moreover, existing cross-view datasets, such as CVUSA [workman2015wide] and VIGOR [zhu2021vigor], were originally designed for cross-view retrieval tasks and thus offer limited diversity for cross-view synthesis.
+To address this, we introduce the Ground2Aerial-3 (G2A-3) dataset, which includes three novel aerial image generation scenarios: disaster emergency response, low-altitude UAV imagery, and historical high-resolution imagery generation tasks.
+The Ground2Aerial-3 dataset provides greater practical value and significance for exploring cross-view synthesis tasks in more diverse scenarios.
 Our main contributions are as follows:
@@
 It also includes "Multi-to-One" mapping strategy to enhance BEV perception range in densely occluded urban areas.
-  • We introduce Ground2Aerial-3, a new ground-to-aerial image synthesis dataset, featuring disaster scene aerial image, historical high-resolution satellite image, and low-altitude UAV image synthesis tasks.
+  • We introduce Ground2Aerial-3, a new ground-to-aerial image synthesis dataset, featuring disaster scene aerial image, low-altitude UAV image, and historical high-resolution satellite image synthesis tasks.
   • SkyDiffusion outperforms state-of-the-art methods on cross-view datasets across natural (CVUSA), suburban (CVACT), urban (VIGOR-Chicago), and various application scenarios (G2A-3), with an SSIM increase of 8.67% and a FID reduction of 21.50%.
-2 Related Work
-Cross-view image synthesis involves generating realistic scenes images with significant viewpoint changes, primarily divided into aerial-to-ground [29, 17, 22] and ground-to-aerial generation [23, 24].
+Related Work
+Cross-view image synthesis involves generating realistic scene images with significant viewpoint changes [shi2022geometry, lu2020geometry, Sat2Density, li2024crossviewdiff, xu2024geospecific].
+This task is common in computer vision and has broad applications in areas such as remote sensing image processing, virtual reality, and autonomous driving.
+It is mainly divided into aerial-to-ground and ground-to-aerial generation [X-Seq, regmi2019cross].
 In this field, synthesizing satellite images from street-views is a challenging yet important task.
-Regmi and Borji [23] attempted to synthesize aerial views from a single street image, but the significant domain differences led to inconsistencies.
-Tang et al.
-[33] used street-view images and aerial semantic maps for image-to-image translation to synthesize target satellite images, but semantic maps are often unavailable in real scenarios.
+X-Seq attempted to synthesize aerial views from a single street image, but the significant domain differences led to inconsistencies.
+tang2019multi used street-view images and aerial semantic maps for image-to-image translation to synthesize target satellite images, but semantic maps are often unavailable in real scenarios.
 Additionally, the performance of GANs has also limited the quality of ground-to-aerial image synthesis in previous studies.
-Concurrent work [1] also synthesizes satellite images using text guidance and staged BEV semantic segmentation results, but these methods require additional text data and an extra semantic segmentation phase on street-view images.
+Concurrent work [arrabi2024cross] also synthesizes satellite images using text guidance and staged BEV semantic segmentation results, but these methods require additional text data and an extra semantic segmentation phase on street-view images.
+Unlike our method, it uses existing BEV methods and only supports one-to-one synthesis.
 Figure 2: Overview of the proposed SkyDiffusion framework.
-It include the curved BEV transformation and BEV-controlled diffusion model with light manipulation.
+It include the curved BEV transformation and BEV-controlled diffusion model.
 The lower parts present the results of One-to-One and Multi-to-One BEV transformations, respectively.
-BEV transformation is a perspective domain conversion technique similar to polar transformation [28, 35] and geo-transformation [17], widely used in retrieval [42], navigation [15] and localization tasks [32, 39, 30].
-Traditional BEV transformation techniques, which rely on camera parameters [27] or depth estimation methods [34, 41], face challenges such as the difficulty in obtaining camera parameters and the high computational cost of depth estimation.
-Recently, some methods have been proposed to map street-view images to the BEV perspective using geometric relations at the image [39] or feature level [30].
+BEV transformation is a perspective domain conversion technique similar to polar transformation [shi2019spatial, toker2021coming] and geo-transformation [lu2020geometry], widely used in retrieval [ye2024cross], navigation [bevformer] and localization tasks [song2024learning, wang2024fine, shi2023boosting].
+Traditional BEV transformation techniques, which rely on camera parameters [sarlin2023orienternet] or depth estimation methods [teng2024360bev, ye2024sg], face challenges such as the difficulty in obtaining camera parameters and the high computational cost of depth estimation.
+Recently, some methods have been proposed to map street-view images to the BEV perspective using geometric relations at the image [wang2024fine] or feature level [shi2023boosting].
 However, such methods are based on the ground plane assumption, where all mapped points are located below the camera, limiting the mapping of upper street scene features and causing information loss.
@@
 Novel view synthesis based on diffusion models.
-Diffusion models have shown significant potential in recent studies on novel view synthesis [7, 4, 37].
-DreamBooth [26] and AerialDiffusion [9] utilize text guidance to achieve top-down image generation by fine-tuning diffusion models.
+Diffusion models have shown significant potential in recent studies on novel view synthesis [kawar2023imagic, gao2023magicdrive, consis].
+DreamBooth [ruiz2023dreambooth] and AerialDiffusion [AerialDiffusion] utilize text guidance to achieve top-down image generation by fine-tuning diffusion models.
 However, it is challenging to fully encapsulate the complexity of real-world scenes using textual information alone.
-Besides, diffusion-based image-to-image translation methods can also generate realistic satellite images [12, 2, 43].
+Besides, diffusion-based image-to-image translation methods can also generate realistic satellite images [li2023bbdm, insd, control].
 However, the substantial domain gap between street-view images and satellite images presents a significant challenge in generating content-consistent satellite images.
-3 Methods
+Methods
 As illustrated in Fig.
 2, this paper introduces SkyDiffusion, a novel method for synthesizing satellite images from corresponding street-view images.
-It initially applies a curved BEV transformation to the input street-view images, converting the perspective to a top-down view using either one-to-one or multi-to-one mappings (Section 3.1).
-Subsequently, a BEV-guided diffusion model controls the synthesis of satellite images, with an optional light manipulation module included(Section 3.2).
-Following this, we detail our network training process (Section 3.3).
-3.1 Curved BEV Transformation
+It initially applies a curved BEV transformation to the input street-view images, converting the perspective to a top-down view using either one-to-one or multi-to-one mappings (Section Curved BEV Transformation).
+Subsequently, a BEV-guided diffusion model controls the synthesis of satellite images (Section BEV-Controlled Diffusion Model).
+Following this, we detail our network training process (Section Network Training and Implementation Details).
+Curved BEV Transformation
 One-to-One BEV.
-Existing BEV methods based on ground plane assumptions ($P(x,y,z=0)$) and geometric relationships achieve the mapping from street-view to the BEV plane[30, 39].
+Existing BEV methods based on ground plane assumptions ($P(x,y,z=0)$) and geometric relationships achieve the mapping from street-view to the BEV plane[shi2023boosting, wang2024fine].
 However, this assumption fails to map the content above the street-view image and results in noticeable distortion as points move away from the center area.
@@
 1 and Eq.
-2, we achieve the projection from the street-view image to the bird’s eye view without requiring camera calibration parameters or depth estimation methods [34, 41].
+2, we achieve the projection from the street-view image to the bird’s eye view without requiring camera calibration parameters or depth estimation methods [teng2024360bev, ye2024sg].
 This effectively maps upper street-view features like distant roads and buildings.
@@
 We select the street-view mapping result closest to the capture point for overlapping areas following Eq.
-3, where $\textit{BEV}\left(x,y\right)$ represents the unified BEV space corresponding to the satellite image, $x,y$ corresponds to the index position of the BEV plane, $\Delta x_{k}$ and $\Delta y_{k}$ represent the offset of the street-view image from the satellite center, and $x_{\mathrm{cam}_{i}}$ and $y_{\mathrm{cam}_{i}}$ represent the coordinates of camera position in the satellite coordinate system.
+3.
 $&\textit{BEV}\left(x,y\right)=\textit{BEV}_{\mathrm{cam}_{k}}\left(x-\Delta x_{k},y-\Delta y_{k}\right)\\ &k=\arg\min_{i}\sqrt{(x-x_{\mathrm{cam}_{i}})^{2}+(y-y_{\mathrm{cam}_{i}})^{2}}\end{aligned}\right.$ (3)
-This enables us to synthesize large-scale scenes using street-view images with limited local information.
-3.2 BEV-Controlled Diffusion Model
-Curved BEV controlled synthesis.
+where $\textit{BEV}\left(x,y\right)$ represents the unified BEV space corresponding to the satellite image, $x,y$ corresponds to the index position of the BEV plane, $\Delta x_{k}$ and $\Delta y_{k}$ represent the offset of the street-view image from the satellite center, and $x_{\mathrm{cam}_{i}}$ and $y_{\mathrm{cam}_{i}}$ represent the coordinates of camera position in the satellite coordinate system.
+Additionally, during the implementation process, we also include boundary cropping as part of the image post-processing.
+BEV-Controlled Diffusion Model
 Image diffusion models progressively denoise images and generate samples from the training domain.
-The denoising process can occur in pixel space or in a latent space encoded from training data.
-Inspired by ControlNet [43], we enhance the control over the generated aerial views by replicating the SD module, enabling the transformed BEV images to effectively guide the synthesis process.
+The denoising process can occur in pixel space or in a latent space encoded from training data [ddim, ddpm].
+In the forward process, diffusion models gradually add Gaussian noises to a ground truth image $x_{0}$ according to a predetermined schedule $\beta_{1},\beta_{2},\dots,\beta_{T}$:
+$q(x_{t}|x_{t-1})=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}I)$ (4)
+where $x_{t}$ is a noised sample with noise level $t$.
+The reverse process involves a series of denoising steps, where noise is progressively removed by employing a neural network $\epsilon_{\theta}$ with parameters $\theta$.
+The reverse process involves a series of denoising steps, where noise is progressively removed by employing a neural network $\epsilon_{\theta}$ with parameters $\theta$.
+This neural network predicts the noise $\epsilon$ present in a noisy image $x_{t}$ at step $t$.
+In the satellite image synthesis process, an effective feature embedding mechanism is required to ensure the generated satellite image aligns with the BEV image $I_{bev}$.
+We use a conditional encoder to transform the BEV image into a latent space feature embedding, and then, similar to ControlNet [control], inject the features into the diffusion model using zero convolution layers and replicated SD modules.
 Specifically, we first use a small encoder to capture the cross-view associative information of the input BEV image $I_{bev}$ such as road directions and building positions, to achieve BEV image feature embedding $c_{bev}$.
@@
 Guided by the BEV image, the diffusion model incorporates cross-view associative information, synthesizing aerial views with consistent content.
-Optional Light Manipulation.
-Given the differences in perspectives and timing between satellite and street images, there is no direct correlation in lighting conditions.
-To address this, we designed a light manipulation module to make the lighting conditions of the synthesized satellite images more adaptable.
-We employ the DNCM for modeling deterministic color mapping [8].
-Initially, we use nDNCM to obtain the "image content information" $I_{c}$ from the preliminary synthesis results $I_{0}$ of the satellite images.
-Then, we input the lighting and color style $r_{s}$ obtained from the reference aerial view into sDNCM for lighting transfer.
-{$&I_{s}=s\textit{DNCM}(\textit{nDNCM}(I_{0},I_{c}),r_{s})\\ &\textit{DNCM}(I,T)=(I_{h\times w,3}\cdot P_{3,k})\cdot T_{k,k}\cdot Q_{k,3}\end{aligned}\right.$ (5)
-$I_{c}$ and $r_{s}$ are the content information and color reference information extracted by the encoder $E$, as shown in Fig.
-2.
-The DNCM module internally includes ${P}_{3,k}$ and ${Q}_{k,3}$, which represent the color embedding and recovery matrices, respectively.
-Here, $k$ represents the embedding dimension.
-The adaptive color mapping matrix ${T}_{k,k}$ is also obtained through the encoder.
-Notably, the lighting adjustment module is independent of the satellite image synthesis, meaning that satellite images can be generated without relying on this module.
-We obtain lighting information from randomly selected satellite images in the training set, and standard images can also be used as lighting control conditions.
-3.3 Network Training
-Loss Function.
-We train the BEV-controlled synthesis of satellite images and the light manipulation module in two separate stages.
-Given a set of cross-view image data, the image diffusion algorithm gradually adds noise to the satellite image, generating a noisy satellite image $Z_{t}$.
+Network Training and Implementation Details
+Given a set of cross-view image data, the image diffusion algorithm gradually adds noise to the satellite image, generating a noisy satellite image.
 It then learns a network $\epsilon_{\theta}$ to predict the noise added to the noisy satellite image $Z_{t}$ based on the given BEV image as the conditional input.
@@
 The loss $\mathcal{L}$ is the learning objective of the diffusion model, where $t$ represents the time step.
-Implementation Details.
-We implement SkyDiffusion using the pre-trained Stable Diffusion v1.5 model [25], with an unlocked diffusion decoder and a classifier-free guidance scale of 9.0 [6].
-Final inference sampling uses 50 steps with the DDIM strategy [31].
+We implement SkyDiffusion using the pre-trained Stable Diffusion v1.5 model [stable], with an unlocked diffusion decoder and a classifier-free guidance scale of 9.0 [cfg].
+Final inference sampling uses 50 steps with the DDIM strategy [ddim].
 Training is on eight NVIDIA A100 GPUs with a batch size of 128 for 100 epochs.
-For optional light manipulation, the reference satellite image is randomly selected from normally toned images in the training set, and the DNCM mapping matrix parameter $k$ is set to 32 (details in Supplementary Material).
-4 Dataset
-We propose Ground2Aerial-3, a multi-task cross-view synthesis dataset designed to explore the performance of cross-view synthesis methods in several novel scenarios.
-As shown in Figure 3, G2A-3 contains nearly 10k pairs of street-view and aerial images, covering disaster scene aerial image synthesis, historical high-resolution satellite image synthesis, and low-altitude UAV image synthesis.
-The dataset of each task is randomly split into training and test sets with a ratio of 5:1.
-The ground street-view images are 1024×512, with true north aligned at the center, and the aerial images are 512×512, aligned with the center of the ground images.
-4.1 Disaster Scene Aerial image
-Both satellite and aerial images are limited by their acquisition frequency, which imposes certain constraints on rapid response to natural disasters.
-Synthesizing aerial images based on street-view images captured by ground cameras can facilitate fast disaster assessment and analysis.
-CVIAN [13] is a cross-view dataset containing street-view and high-resolution satellite imagery in Florida, USA, following Hurricane Ian in 2022.
-From CVIAN’s street-view and satellite data, we filtered 2.7k image pairs of moderately or severely affected areas and adjusted the north direction of street-view images to the central column, preparing the data for the disaster scene aerial image generation task.
-Figure 4: Illustration of the Ground2Aerial-3 dataset.
-4.2 Historical High-Resolution Satellite image
-Obtaining high-resolution satellite imagery, particularly ultra-high resolution images exceeding 0.3m, has been challenging.
-It wasn’t until the recent launch of WorldView-3 (2014) and WorldView-4 (2018) that high-resolution coverage at near-global scales became available.
-Previously, ultra-high resolution images were often limited to small areas captured by aerial photography, with both spatial and temporal coverage constraints.
-Synthesizing satellite images based on historical street-view images can help fill this gap, providing additional complementary information.
-We collected historical street-view images from 2007 to 2014 on the Mapillary  [^1] https://www.mapillary.com/ platform and used corresponding aerial imagery from the massGIS platform  [^2] https://www.mass.gov/info-details/massgis-data-layers as ground truth to conduct high-resolution satellite image synthesis tasks.
-4.3 Low-Altitude UAV Image
-Low-altitude UAV imagery, with ultra-high resolution below 0.05 meters, has potential applications in tasks such as lane data collection for autonomous driving.
-However, the coverage and cost of UAV imagery are limited, and safety regulations further restrict coverage range and update frequency.
-Synthesizing low-altitude UAV images from more accessible street-view images presents a potential application scenario.
-Using virtual MatrixCity [14] data, we employed the UE engine to set up six single-view cameras at ground level, rendering and stitching them into panoramic street-view images.
-At the same locations, we positioned downward-facing cameras at an altitude of 20 meters to render low-altitude UAV imagery, creating a cross-view dataset for low-altitude UAV image synthesis tasks.
-Figure 5: Qualitative comparison of different methods synthesis results on three datasets.
-5 Experiments
-In this section, we introduce the datasets utilized and the experimental setup.
-We then conduct both qualitative and quantitative comparisons of SkyDiffusion with state-of-the-art cross-view synthesis methods.
-Finally, we perform ablation studies to evaluate the effectiveness of each module.
-5.1 Experimental Setup
-Dataset Details.
-CVUSA [40] and CVACT [16] are two commonly used cross-view datasets, containing one-to-one, center-aligned aerial and ground street-view images.
-They mainly covers suburban scenes, with low-rise buildings and open views.
-The CVACT dataset is regionally segmented.
-Although CVUSA uses random partitioning, its data distribution is sparse, with no overlap among satellite images.
-VIGOR-Chicago [46] is a subset of the VIGOR cross-view dataset, containing the Chicago area.
-In this dataset, street-view panoramas and satellite images are not center-aligned, with multiple street-view images covering the same satellite image region, forming a multi-to-one mapping relationship.
-We modified the train-test split to a 4:1 ratio and ensured that the same satellite images do not appear in both the training and testing sets to prevent information leakage.
-The G2A-3 dataset has been described in Section 4.
-All input street images are set to 512 $\times$ 1024 pixels, with the center column aligned to the north.
-Unlike previous methods that synthesize low-resolution satellite images (64 $\times$ 64), this experiment aims to meet practical needs by generating higher-resolution satellite images (512 $\times$ 512 pixels).
-In the Curved-BEV transformation, $\lambda$ is set to 2 and $H$ to 1.5 for CVUSA, CVACT and G2A-3 datasets.
-For the VIGOR-Chicago dataset, when performing a one-to-one mapping, $\lambda$ is set to 10; for a Multi-to-One mapping, $\lambda$ is set to 0, the primary focus is on retaining the street-view central location information.
-The Curved-BEV transformation generates results that are size-matched with satellite images.
-Comparison Methods and Metric.
-We compared our method with several state-of-the-art cross-view synthesis methods, including the commonly used cross-view satellite synthesis methods X-Seq [23], SelectionGAN (without semantic masks) [33], and CDTE [35], image translation methods such as the GAN-based CUT [20] and diffusion-based Instruct pix2pix (Instr-p2p) [2], ControlNet [43] and img2img-turbo [21], as well as novel view synthesis methods like Aerial Diff [9] and HawkI [10].
-Except for the fine-tuned diffusion model, which requires brief textual descriptions, all other models use the same street-view input.
-For a fair comparison, the results in the table 1 are presented without the use of the Optional Light Manipulation module for Skydiffusion.
-Following previous work [17, 35, 23], we employ commonly used metrics such as SSIM, PSNR and LPIPS to evaluate the content consistency of the synthesized images, and use FID [5] to assess the realism of images.
-Method CVUSA CVACT VIGOR-Chicago
-FID ($\downarrow$) SSIM ($\uparrow$) PSNR ($\uparrow$) LPIPS ($\downarrow$) FID ($\downarrow$) SSIM ($\uparrow$) PSNR ($\uparrow$) LPIPS ($\downarrow$) FID ($\downarrow$) SSIM ($\uparrow$) PSNR ($\uparrow$) LPIPS ($\downarrow$)
-X-Seq [23] 161.16 0.084 11.97 0.706 190.12 0.042 12.41 0.661 - - - -
-CDTE [35] 122.84 0.143 10.01 0.694 160.81 0.091 12.59 0.663 - - - -
-SelGAN [33] 116.57 0.129 12.38 0.742 100.21 0.116 12.04 0.684 149.53 0.127 11.77 0.778
-Aerial Diff [9] 136.18 0.103 10.06 0.855 127.29 0.108 10.24 0.878 123.16 0.141 11.49 0.831
-HawkI [10] 90.98 0.114 11.09 0.783 135.75 0.112 10.97 0.764 123.47 0.124 9.67 0.752
-CUT [20] 72.83 0.121 12.09 0.687 62.22 0.102 13.11 0.664 69.42 0.169 13.66 0.665
-I2I-Turbo [21] 77.95 0.127 12.64 0.685 73.24 0.103 12.19 0.679 80.10 0.135 11.28 0.690
-Instr-p2p [2] 38.01 0.138 12.29 0.697 49.62 0.116 12.51 0.682 50.30 0.163 10.62 0.689
-ControlNet [43] 32.45 0.149 12.63 0.650 62.21 0.115 11.95 0.682 53.27 0.170 10.38 0.666
-Ours 29.18 0.168 14.58 0.635 36.48 0.118 12.85 0.645 45.29 0.186 11.69 0.661
-Table 1: Quantitative comparison of different methods on CVUSA, CVACT and VIGOR-Chicago.
-Method FID ($\downarrow$) SSIM ($\uparrow$) PSNR ($\uparrow$) LPIPS ($\downarrow$)
-Disaster Scene image synthesis
-ControlNet 138.11 0.183 12.95 0.690
-Ours 114.64 0.188 13.02 0.694
-Historical aerial image synthesis
-ControlNet 36.48 0.117 12.85 0.665
-Ours 35.51 0.123 13.58 0.657
-Low-altitude UAV image synthesis
-ControlNet 52.76 0.371 16.55 0.598
-Ours 32.11 0.452 16.86 0.390
-Table 2: Quantitative comparison of methods on G2A-3 Dataset.
-5.2 Comparison with State-of-the-Art Methods
-On the suburban CVUSA and CVACT datasets, SkyDiffusion achieved the outstanding results, as shown in Table 1.
-Compared to state-of-the-art methods, it reduced FID by 25.72% and increased SSIM by 7.68%, demonstrating its superiority in synthesizing realistic and consistent satellite images.
-The visual results in Fig.
-5 show that GAN-based cross-view methods generate excessive artifacts and blurriness.
-While diffusion-based methods such as Aerial Diff and Instr-p2p produce highly realistic aerial views, they lack content correlation with the street-view images.
-In contrast, our method benefits from BEV perspective transformation, enabling effective capture of cross-view associative information.
-SkyDiffusion generates realistic images with consistent content layout, including road direction, tree placement, and building distribution.
-Additional visual results can be found in the supplementary materials.
-Figure 6: Qualitative Visualization Results on the G2A-3 Dataset.
-In the urban VIGOR-Chicago dataset, our method reduced FID by 14.98% and improved SSIM by 9.41% compared to the state-of-the-art method, as shown in Table 1.
-The VIGOR results for SkyDiffusion use only a single street-view image.
-The results using Multi-to-One BEV generation with multiple street-view images can be found in the ablation study.
-The visual results in Fig.
-5 show that our method maintains consistent associative information such as crosswalks and lanes.
-While the non-associative information on building rooftops cannot be consistently generated, it still produces reasonable and realistic results.
-Figure 7: Qualitative comparison of synthesis results for One-to-One and Multi-to-one BEV.
-As shown in Table 2, the tasks on the G2A-3 dataset present certain challenges; however, our method achieves significant performance improvements over the commonly used image-conditioned synthesis method, ControlNet.
-SkyDiffusion reduces the FID by an average of 19.60% and increases the SSIM by an average of 9.90%.
-The visual results in Figure 6 demonstrate SkyDiffusion’s ability to synthesize various types of aerial images across different G2A-3 task settings.
-In Figure 6(a), prominent disaster areas in the street-view image (orange box indicating debris of damaged buildings and red box showing trees downed by the hurricane) are clearly visible in the synthesized satellite image, aiding in subsequent disaster area assessment and analysis.
-Figure 6(b) shows that SkyDiffusion can synthesize realistic early historical imagery, effectively filling gaps where early high-resolution aerial images are lacking.
-Figure 6(c) highlights SkyDiffusion’s exceptional performance in low-altitude UAV image synthesis.
-Due to the lower altitude of UAV captures, there is greater overlap and closer domain similarity between scenes, resulting in superior synthesis quality.
-SkyDiffusion demonstrates excellent aerial image synthesis capabilities, advancing potential applications in various aerial image synthesis scenarios.
-5.3 Ablation Study and More Exploration
-Curved-BEV controlled synthesis.
-We conducted ablation experiments on the three datasets, with results shown in Table 3.
-Compared to directly using street-view images as input, the Curved-BEV method improves performance across multiple metrics by transforming street-view images into satellite views for domain alignment.
-This indicates that the Curved-BEV method aids in synthesizing more content-consistent satellite images.
-Furthermore, the Multi-to-one method further improves metrics compared to the one-to-one mapping, demonstrating its effectiveness in dense urban scenes.
-As shown in the visual results in Figure 7, using multiple street-view images for multi-to-one BEV mapping effectively addresses occlusions from tall objects like trees and buildings.
-This method effectively enhances the BEV sensing range in urban environments, facilitating the generation of large-scale satellite scenes.
-Additional ablation study results can be found in the supplementary materials.
-C-BEV Light FID ($\downarrow$) SSIM ($\uparrow$) PSNR ($\uparrow$) LPIPS ($\downarrow$)
-CVACT 62.21 0.115 11.95 0.682
-36.48 0.118 12.85 0.645
-35.51 0.123 13.58 0.657
-VIGOR 53.27 0.170 10.38 0.666
-45.29 0.186 11.69 0.661
-(Multi) 37.90 0.193 10.39 0.660
-(Multi) 31.90 0.205 11.15 0.651
-Table 3: Ablation study of the Curved-BEV and optional Light Manipulation Modules. "Multi" stands for Multi-to-One strategy, and "Light" represents using light manipulation.
-Light manipulation.
-Aerial and satellite images typically capture large areas with relatively uniform overall brightness.
-Using small aerial image patches as references effectively captures the lighting conditions at the time of capture.
-We applied an optional Light Manipulation module to align the lighting conditions of the synthesized image with those of the target domain image, as shown in Table 3 The results indicate that Light Manipulation can effectively improve SSIM and PSNR metrics.
-This module preserves the original image content while providing more flexible lighting conditions for the synthesized satellite image.
-Additional visualizations can be found in the supplementary material.
-6 Conclusion
-In this study, we introduce SkyDiffusion, a novel approach specifically designed for street images to satellite images cross-view synthesis.
-SkyDiffusion operates solely with street images as input, utilizing a BEV Paradigm and diffusion models to generate satellite images.
-SkyDiffusion achieves state-of-the-art performance in both content consistency and image realism on across multiple cross-view datasets, demonstrating its superior capabilities.
-Additionally, we introduce a cross-view synthesis dataset, Ground2Aerial-3, featuring aerial image synthesis tasks for multiple new scenes, providing practical value and inspiration for future cross-view synthesis research.

### v5

@@
-1Association for the Advancement of Artificial Intelligence
+Leveraging BEV Paradigm for Ground-to-Aerial Image Synthesis
 Abstract
-Street-to-satellite image synthesis focuses on generating realistic satellite images from corresponding ground street-view images while maintaining a consistent content layout, similar to looking down from the sky.
-The significant differences in perspectives create a substantial domain gap between the views, making this cross-view generation task particularly challenging.
-In this paper, we introduce SkyDiffusion, a novel cross-view generation method for synthesizing satellite images from street-view images, leveraging diffusion models and Bird’s Eye View (BEV) paradigm.
-First, we design a Curved-BEV method to transform street-view images to the satellite view, reformulating the challenging cross-domain image synthesis task into a conditional generation problem.
-Curved-BEV also includes a "Multi-to-One" mapping strategy for leveraging multiple street-view images within the same satellite coverage area, effectively solving the occlusion issues in dense urban scenes.
-Next, we design a BEV-controlled diffusion model to generate satellite images consistent with the street-view content, which also incorporates a light manipulation module to make the lighting conditions of the synthesized satellite images more flexible.
-Experimental results demonstrate that SkyDiffusion outperforms state-of-the-art methods on both suburban (CVUSA & CVACT) and urban (VIGOR-Chicago) cross-view datasets, with an average SSIM increase of 13.96% and a FID reduction of 20.54%, achieving realistic and content-consistent satellite image generation.
-The code and models will be made publicly available.
-Introduction
-In fields such as land cover classification [lv2023novel, tong2020land], urban planning [zhu2019understanding, zhao2019exploring], and disaster response [kucharczyk2021remote, zhao2024see, wang2023hierarchical], the analysis of aerial imagery plays a crucial role.
+Ground-to-aerial image synthesis focuses on generating realistic aerial images from corresponding ground street view images while maintaining consistent content layout, simulating a top-down view.
+The significant viewpoint difference leads to domain gaps between views, and dense urban scenes limit the visible range of street views, making this cross-view generation task particularly challenging.
+In this paper, we introduce SkyDiffusion, a novel cross-view generation method for synthesizing aerial images from street view images, utilizing a diffusion model and the Bird’s-Eye View (BEV) paradigm.
+The Curved-BEV method in SkyDiffusion converts street-view images into a BEV perspective, effectively bridging the domain gap, and employs a "multi-to-one" mapping strategy to address occlusion issues in dense urban scenes.
+Next, SkyDiffusion designed a BEV-guided diffusion model to generate content-consistent and realistic aerial images.
+Additionally, we introduce a novel dataset, Ground2Aerial-3, designed for diverse ground-to-aerial image synthesis applications, including disaster scene aerial synthesis, low-altitude UAV image synthesis, and historical high-resolution satellite image synthesis tasks.
+Experimental results demonstrate that SkyDiffusion outperforms state-of-the-art methods on cross-view datasets across natural (CVUSA), suburban (CVACT), urban (VIGOR-Chicago), and various application scenarios (G2A-3), achieving realistic and content-consistent aerial image generation.
+The code, datasets and more information of this work can be found at https://opendatalab.github.io/skydiffusion/.
+1 Introduction
+In fields such as land cover classification [21, 38], urban planning [56, 53, 54], and disaster response [10, 52, 40], the analysis of aerial imagery plays a crucial role.
 The cross-view Ground-to-Aerial Image Synthesis task leverages ground street-view images to synthesize corresponding aerial images at the same location.
-Street-view images can be captured and uploaded by users on platforms such as Google Street View or Mapillary, offering high flexibility in data collection and frequent temporal coverage [arrabi2024cross].
-Therefore, Ground-to-Aerial Image Synthesis can support rapid response during natural disasters [ou2023method] or fill historical gaps in high-resolution satellite imagery records [cornebise2022open], among other applications.
+Street-view images can be captured and uploaded by users on platforms such as Google Street View or Mapillary, offering high flexibility in data collection and frequent temporal coverage [1].
+Therefore, Ground-to-Aerial Image Synthesis can support rapid response during natural disasters [22] or fill historical gaps in high-resolution satellite imagery records [3], among other applications.
 Figure 1: Illustration of the cross-view image synthesis task.
@@
 Despite the potential of cross-view synthesis from ground-level to aerial imagery, significant challenges remain.
-The drastic viewpoint differences between street and satellite views lead to distinct object appearances, resulting in substantial domain gaps across image types.
+The drastic viewpoint differences between street and satellite views lead to distinct object appearances, resulting in substantial domain gaps across image types[18, 13].
 As shown in Figure 1 (a), in cross-view scenarios, street-view images primarily capture ground and building facade details, while satellite views present more rooftop and macroscopic summary information.
 Even with state-of-the-art diffusion methods, the domain gap between views often results in generated images that are realistic yet not true to reality; the synthesized satellite imagery may differ significantly from the actual content of corresponding street-views, as shown in Figure 1(c).
-Another challenge is object occlusion in dense urban settings, where tall buildings or trees limit the visibility range in street-views, which is substantially smaller than the coverage of satellite images, as shown in Figure 1(b).
+Another challenge is object occlusion in dense urban settings, where tall buildings or trees limit the visibility range in street-views, which is substantially smaller than the coverage of satellite images [15], as shown in Figure 1(b).
 Aerial images synthesized from a single street-view image struggle to maintain consistency with real satellite images in unseen areas.
@@
 Therefore, SkyDiffusion effectively overcomes the challenges of cross-view domain gaps and the limited observation range in densely occluded urban scenes, generating more realistic satellite images.
-Moreover, existing cross-view datasets, such as CVUSA [workman2015wide] and VIGOR [zhu2021vigor], were originally designed for cross-view retrieval tasks and thus offer limited diversity for cross-view synthesis.
+Moreover, existing cross-view datasets, such as CVUSA [42] and VIGOR [55], were originally designed for cross-view retrieval tasks and thus offer limited diversity for cross-view synthesis.
 To address this, we introduce the Ground2Aerial-3 (G2A-3) dataset, which includes three novel aerial image generation scenarios: disaster emergency response, low-altitude UAV imagery, and historical high-resolution imagery generation tasks.
@@
   • SkyDiffusion outperforms state-of-the-art methods on cross-view datasets across natural (CVUSA), suburban (CVACT), urban (VIGOR-Chicago), and various application scenarios (G2A-3), with an SSIM increase of 8.67% and a FID reduction of 21.50%.
-Related Work
-Cross-view image synthesis involves generating realistic scene images with significant viewpoint changes [shi2022geometry, lu2020geometry, Sat2Density, li2024crossviewdiff, xu2024geospecific].
+2 Related Work
+Cross-view image synthesis involves generating realistic scene images with significant viewpoint changes [49, 25, 14, 43, 51, 48].
 This task is common in computer vision and has broad applications in areas such as remote sensing image processing, virtual reality, and autonomous driving.
-It is mainly divided into aerial-to-ground and ground-to-aerial generation [X-Seq, regmi2019cross].
+It is mainly divided into aerial-to-ground and ground-to-aerial generation [26, 27].
 In this field, synthesizing satellite images from street-views is a challenging yet important task.
-X-Seq attempted to synthesize aerial views from a single street image, but the significant domain differences led to inconsistencies.
-tang2019multi used street-view images and aerial semantic maps for image-to-image translation to synthesize target satellite images, but semantic maps are often unavailable in real scenarios.
+Regmi and Borji [26] attempted to synthesize aerial views from a single street image, but the significant domain differences led to inconsistencies.
+Tang et al.
+[35] used street-view images and aerial semantic maps for image-to-image translation to synthesize target satellite images, but semantic maps are often unavailable in real scenarios.
 Additionally, the performance of GANs has also limited the quality of ground-to-aerial image synthesis in previous studies.
-Concurrent work [arrabi2024cross] also synthesizes satellite images using text guidance and staged BEV semantic segmentation results, but these methods require additional text data and an extra semantic segmentation phase on street-view images.
+Concurrent work [1] also synthesizes satellite images using text guidance and staged BEV semantic segmentation results, but these methods require additional text data and an extra semantic segmentation phase on street-view images.
 Unlike our method, it uses existing BEV methods and only supports one-to-one synthesis.
@@
 The lower parts present the results of One-to-One and Multi-to-One BEV transformations, respectively.
-BEV transformation is a perspective domain conversion technique similar to polar transformation [shi2019spatial, toker2021coming] and geo-transformation [lu2020geometry], widely used in retrieval [ye2024cross], navigation [bevformer] and localization tasks [song2024learning, wang2024fine, shi2023boosting].
-Traditional BEV transformation techniques, which rely on camera parameters [sarlin2023orienternet] or depth estimation methods [teng2024360bev, ye2024sg], face challenges such as the difficulty in obtaining camera parameters and the high computational cost of depth estimation.
-Recently, some methods have been proposed to map street-view images to the BEV perspective using geometric relations at the image [wang2024fine] or feature level [shi2023boosting].
+BEV transformation is a perspective domain conversion technique similar to polar transformation [31, 37] and geo-transformation [20], widely used in retrieval [46], navigation [17] and localization tasks [34, 41, 32].
+Traditional BEV transformation techniques, which rely on camera parameters [30] or depth estimation methods [36, 45], face challenges such as the difficulty in obtaining camera parameters and the high computational cost of depth estimation.
+Recently, some methods have been proposed to map street-view images to the BEV perspective using geometric relations at the image [41] or feature level [32].
 However, such methods are based on the ground plane assumption, where all mapped points are located below the camera, limiting the mapping of upper street scene features and causing information loss.
@@
 Novel view synthesis based on diffusion models.
-Diffusion models have shown significant potential in recent studies on novel view synthesis [kawar2023imagic, gao2023magicdrive, consis].
-DreamBooth [ruiz2023dreambooth] and AerialDiffusion [AerialDiffusion] utilize text guidance to achieve top-down image generation by fine-tuning diffusion models.
+Diffusion models have shown significant potential in recent studies on novel view synthesis [44, 8, 4, 47, 39].
+DreamBooth [29] and AerialDiffusion [9] utilize text guidance to achieve top-down image generation by fine-tuning diffusion models.
 However, it is challenging to fully encapsulate the complexity of real-world scenes using textual information alone.
-Besides, diffusion-based image-to-image translation methods can also generate realistic satellite images [li2023bbdm, insd, control].
+Besides, diffusion-based image-to-image translation methods can also generate realistic satellite images [11, 2, 50].
 However, the substantial domain gap between street-view images and satellite images presents a significant challenge in generating content-consistent satellite images.
-Methods
+3 Methods
 As illustrated in Fig.
 2, this paper introduces SkyDiffusion, a novel method for synthesizing satellite images from corresponding street-view images.
-It initially applies a curved BEV transformation to the input street-view images, converting the perspective to a top-down view using either one-to-one or multi-to-one mappings (Section Curved BEV Transformation).
-Subsequently, a BEV-guided diffusion model controls the synthesis of satellite images (Section BEV-Controlled Diffusion Model).
-Following this, we detail our network training process (Section Network Training and Implementation Details).
-Curved BEV Transformation
+It initially applies a curved BEV transformation to the input street-view images, converting the perspective to a top-down view using either one-to-one or multi-to-one mappings (Section 3.1).
+Subsequently, a BEV-guided diffusion model controls the synthesis of satellite images (Section 3.2).
+Following this, we detail our network training process (Section 3.3).
+3.1 Curved BEV Transformation
 One-to-One BEV.
-Existing BEV methods based on ground plane assumptions ($P(x,y,z=0)$) and geometric relationships achieve the mapping from street-view to the BEV plane[shi2023boosting, wang2024fine].
+Existing BEV methods based on ground plane assumptions ($P(x,y,z=0)$) and geometric relationships achieve the mapping from street-view to the BEV plane[32, 41].
 However, this assumption fails to map the content above the street-view image and results in noticeable distortion as points move away from the center area.
@@
 More details are in the supplementary material.
-{$u&=\left[\text{arctan2}\left(y,x\right)+\pi\right]\frac{w}{2\pi}\\ v&=\left[\frac{\pi}{2}+\text{arctan2}\left(z-H,\sqrt{x^{2}+y^{2}}\right)\right]\frac{h}{\pi}\end{aligned}\right.$ (2)
+$\left\{\begin{aligned} u&=\left[\text{arctan2}\left(y,x\right)+\pi\right]\frac{w}{2\pi}\\ v&=\left[\frac{\pi}{2}+\text{arctan2}\left(z-H,\sqrt{x^{2}+y^{2}}\right)\right]\frac{h}{\pi}\end{aligned}\right$.
+(2)
 Following Eq.
 1 and Eq.
-2, we achieve the projection from the street-view image to the bird’s eye view without requiring camera calibration parameters or depth estimation methods [teng2024360bev, ye2024sg].
+2, we achieve the projection from the street-view image to the bird’s eye view without requiring camera calibration parameters or depth estimation methods [36, 45].
 This effectively maps upper street-view features like distant roads and buildings.
@@
 3.
-$&\textit{BEV}\left(x,y\right)=\textit{BEV}_{\mathrm{cam}_{k}}\left(x-\Delta x_{k},y-\Delta y_{k}\right)\\ &k=\arg\min_{i}\sqrt{(x-x_{\mathrm{cam}_{i}})^{2}+(y-y_{\mathrm{cam}_{i}})^{2}}\end{aligned}\right.$ (3)
+$\left\{\begin{aligned} &\textit{BEV}\left(x,y\right)=\textit{BEV}_{\mathrm{cam}_{k}}\left(x-\Delta x_{k},y-\Delta y_{k}\right)\\ &k=\arg\min_{i}\sqrt{(x-x_{\mathrm{cam}_{i}})^{2}+(y-y_{\mathrm{cam}_{i}})^{2}}\end{aligned}\right$.
+(3)
 where $\textit{BEV}\left(x,y\right)$ represents the unified BEV space corresponding to the satellite image, $x,y$ corresponds to the index position of the BEV plane, $\Delta x_{k}$ and $\Delta y_{k}$ represent the offset of the street-view image from the satellite center, and $x_{\mathrm{cam}_{i}}$ and $y_{\mathrm{cam}_{i}}$ represent the coordinates of camera position in the satellite coordinate system.
 Additionally, during the implementation process, we also include boundary cropping as part of the image post-processing.
-BEV-Controlled Diffusion Model
+3.2 BEV-Controlled Diffusion Model
 Image diffusion models progressively denoise images and generate samples from the training domain.
-The denoising process can occur in pixel space or in a latent space encoded from training data [ddim, ddpm].
+The denoising process can occur in pixel space or in a latent space encoded from training data [33, 7].
 In the forward process, diffusion models gradually add Gaussian noises to a ground truth image $x_{0}$ according to a predetermined schedule $\beta_{1},\beta_{2},\dots,\beta_{T}$:
@@
 In the satellite image synthesis process, an effective feature embedding mechanism is required to ensure the generated satellite image aligns with the BEV image $I_{bev}$.
-We use a conditional encoder to transform the BEV image into a latent space feature embedding, and then, similar to ControlNet [control], inject the features into the diffusion model using zero convolution layers and replicated SD modules.
+We use a conditional encoder to transform the BEV image into a latent space feature embedding, and then, similar to ControlNet [50], inject the features into the diffusion model using zero convolution layers and replicated SD modules.
 Specifically, we first use a small encoder to capture the cross-view associative information of the input BEV image $I_{bev}$ such as road directions and building positions, to achieve BEV image feature embedding $c_{bev}$.
@@
 Guided by the BEV image, the diffusion model incorporates cross-view associative information, synthesizing aerial views with consistent content.
-Network Training and Implementation Details
+3.3 Network Training and Implementation Details
 Given a set of cross-view image data, the image diffusion algorithm gradually adds noise to the satellite image, generating a noisy satellite image.
@@
 The loss $\mathcal{L}$ is the learning objective of the diffusion model, where $t$ represents the time step.
-We implement SkyDiffusion using the pre-trained Stable Diffusion v1.5 model [stable], with an unlocked diffusion decoder and a classifier-free guidance scale of 9.0 [cfg].
-Final inference sampling uses 50 steps with the DDIM strategy [ddim].
+We implement SkyDiffusion using the pre-trained Stable Diffusion v1.5 model [28], with an unlocked diffusion decoder and a classifier-free guidance scale of 9.0 [6].
+Final inference sampling uses 50 steps with the DDIM strategy [33].
 Training is on eight NVIDIA A100 GPUs with a batch size of 128 for 100 epochs.
+4 Dataset
+We propose Ground2Aerial-3, a multi-task cross-view synthesis dataset designed to explore the performance of cross-view synthesis methods in several novel scenarios.
+As shown in Figure 4, G2A-3 contains around 20,000 street view and aerial images, covering disaster scene aerial image synthesis, low-altitude UAV image synthesis, and historical high-resolution satellite image synthesis.
+Due to the challenges in collecting data for specialized task scenarios, G2A-3 has gathered the available data to the greatest extent possible.
+The ground street-view images are 1024 $\times$ 512, with true north aligned at the center, and the aerial images are 512 $\times$ 512, aligned with the center of the ground images.
+The dataset for each task is divided into training and testing sets by region, with a 4:1 ratio.
+The motivation and specific information for each task is as follows.
+4.1 Disaster Response Image Synthesis
+Satellite and aerial imagery are often limited by their capture frequency, making it difficult to obtain timely data for disaster response scenarios.
+Furthermore, acquiring multiple satellite images within a short time frame (e.g., 1 hour) to track disaster changes is typically infeasible.
+In such cases, synthesizing aerial images from street view images captured by ground cameras offers a fast and cost-effective solution, which can be further used for disaster assessment via comparative analysis with pre-disaster aerial imagery and assist in disaster geo-localization.
+CVIAN [12] is a cross-view dataset containing street-view and high-resolution satellite imagery in Florida, USA, following Hurricane Ian in 2022.
+From CVIAN’s street-view and satellite data, we filtered 2.7k image pairs of moderately or severely affected areas and adjusted the north direction of street-view images to the central column, preparing the data for the disaster scene aerial image generation task.
+4.2 Low-Altitude UAV Image Synthesis
+Low-altitude UAV imagery, with ultra-high resolution below 0.05 meters, has potential applications in tasks such as lane data collection for autonomous driving.
+However, the coverage and cost of UAV imagery are limited, and safety regulations further restrict coverage range and update frequency.
+Synthesizing low-altitude UAV images from more accessible street-view images presents a potential application scenario.
+Using virtual MatrixCity [16] data, we employed the UE engine to set up six single-view cameras at ground level, rendering and stitching them into panoramic street-view images.
+Specific details are provided in the supplementary materials.
+At the same locations, we positioned downward-facing cameras at an altitude of 20 meters to render low-altitude UAV imagery, creating a cross-view dataset for low-altitude UAV image synthesis tasks.
+4.3 Historical HR Satellite Image Synthesis
+Obtaining high-resolution (HR) satellite imagery, particularly ultra-high resolution images exceeding 0.3m, has been challenging.
+It wasn’t until the recent launch of WorldView-3 (2014) and WorldView-4 (2018) that high-resolution coverage at near-global scales became available.
+Previously, ultra-high resolution images were often limited to small areas captured by aerial photography, with both spatial and temporal coverage constraints.
+Therefore, by utilizing historical street view images, we can provide information about past scenes to synthesize satellite or aerial images of the same location, offering additional supplementary information.
+We collected historical street view images from 2007 to 2014 for the Boston and Los Angeles areas on the Mapillary  [^1] https://www.mapillary.com/ and used corresponding aerial images from the massGIS platform  [^2] https://www.mass.gov/info-details/massgis-data-layers as high-resolution satellite imagery for training and validation of synthesis results.
+Due to significant variations in the number of street-view images across different years, we still organized the dataset by region.
+Figure 4: Illustration of the Ground2Aerial-3 dataset.
+Figure 5: Qualitative comparison of different methods synthesis results on three datasets.
+5 Experiments
+In this section, we introduce the datasets utilized and the experimental setup.
+We then conduct both qualitative and quantitative comparisons of SkyDiffusion with state-of-the-art cross-view synthesis methods.
+Finally, we perform ablation studies to evaluate the effectiveness of each module.
+5.1 Experimental Setup
+Dataset Details.
+CVUSA [42] and CVACT [19] are two commonly used cross-view datasets, containing one-to-one, center-aligned aerial and ground street-view images.
+They mainly covers suburban scenes, with low-rise buildings and open views.
+The CVACT dataset is regionally segmented.
+Although CVUSA uses random partitioning, its data distribution is sparse, with no overlap among satellite images.
+VIGOR-Chicago [55] is a subset of the VIGOR cross-view dataset, containing the Chicago area.
+In this dataset, street-view panoramas and satellite images are not center-aligned, with multiple street-view images covering the same satellite image region, forming a multi-to-one mapping relationship.
+We modified the train-test split to a 4:1 ratio and ensured that the same satellite images do not appear in both the training and testing sets to prevent information leakage.
+The G2A-3 dataset has been described in Section 4.
+Parameter setting details.
+All input street images are set to 512 $\times$ 1024 pixels, with the center column aligned to the north.
+Unlike previous methods that synthesize low-resolution satellite images (64 $\times$ 64), this experiment aims to meet practical needs by generating higher-resolution satellite images (512 $\times$ 512 pixels).
+For a multi-to-one mapping, to facilitate the merging of BEV mapping results from different street views, we set the hyperparameter $\lambda$ to 0, allowing for direct nearest-distance integration on the same horizontal plane.
+In this case, the BEV mapping of a single street-view image is similar to those in [32, 46], as it only needs to capture details near the capture point during the multi-to-one mapping process.
+Comparison Methods and Metric.
+We compare our method with several state-of-the-art cross-view synthesis techniques, including GAN-based methods such as X-Seq [26], SelectionGAN (without semantic masks) [35], CDTE [37], and CUT [23], as well as diffusion-based new view synthesis and image transformation methods, including Aerial Diff [9], GPG2A [1], Instruct Pix2Pix (Instr-p2p) [2], ControlNet [50], and img2img-turbo [24].
+Except for the fine-tuned diffusion model, which requires brief textual descriptions, all other models use the same street-view input.
+Following previous work [20, 37, 26], we employ commonly used metrics such as SSIM, PSNR and LPIPS to evaluate the content consistency of the synthesized images, and use FID [5] to assess the realism of images.
+Method CVUSA CVACT VIGOR-Chicago
+FID ($\downarrow$) SSIM ($\uparrow$) PSNR ($\uparrow$) LPIPS ($\downarrow$) FID ($\downarrow$) SSIM ($\uparrow$) PSNR ($\uparrow$) LPIPS ($\downarrow$) FID ($\downarrow$) SSIM ($\uparrow$) PSNR ($\uparrow$) LPIPS ($\downarrow$)
+X-Seq [26] 161.16 0.084 11.97 0.706 190.12 0.042 12.41 0.661 - - - -
+CDTE [37] 122.84 0.143 10.01 0.694 160.81 0.091 12.59 0.663 - - - -
+SelGAN [35] 116.57 0.129 12.38 0.742 100.21 0.116 12.04 0.684 149.53 0.127 11.77 0.778
+Aerial Diff [9] 136.18 0.103 10.06 0.855 127.29 0.108 10.24 0.878 123.16 0.141 11.49 0.831
+CUT [23] 72.83 0.121 12.09 0.687 62.22 0.102 13.11 0.664 69.42 0.169 13.66 0.665
+I2I-Turbo [24] 77.95 0.127 12.64 0.685 73.24 0.103 12.19 0.679 80.10 0.135 11.28 0.690
+GPG2A [1] 58.80 0.135 12.13 0.691 63.50 0.116 11.98 0.690 70.19 0.159 11.81 0.695
+Instr-p2p [2] 38.01 0.138 12.29 0.697 49.62 0.116 12.51 0.682 50.30 0.163 10.62 0.689
+ControlNet [50] 32.45 0.149 12.63 0.650 62.21 0.115 11.95 0.682 53.27 0.170 10.38 0.666
+Ours 29.18 0.168 14.58 0.635 36.48 0.118 12.85 0.645 45.29 0.186 11.69 0.661
+Table 1: Quantitative comparison of different methods on CVUSA, CVACT and VIGOR-Chicago.
+Figure 6: Qualitative comparison of synthesis results for One-to-One and Multi-to-one BEV.
+5.2 Comparison with State-of-the-Art Methods
+On the suburban CVUSA and CVACT datasets, SkyDiffusion achieved the outstanding results, as shown in Table 1.
+Compared to state-of-the-art methods, it reduced FID by 25.72% and increased SSIM by 7.68%, demonstrating its superiority in synthesizing realistic and consistent satellite images.
+The visual results in Fig.
+5 show that GAN-based cross-view methods generate excessive artifacts and blurriness.
+While diffusion-based methods such as Aerial Diff and Instr-p2p produce highly realistic aerial views, they lack content correlation with the street-view images.
+In contrast, our method benefits from BEV perspective transformation, enabling effective capture of cross-view associative information.
+SkyDiffusion generates realistic images with consistent content layout, including road direction, tree placement, and building distribution.
+Additional visual results can be found in the supplementary materials.
+In the urban VIGOR-Chicago dataset, our method reduced FID by 14.98% and improved SSIM by 9.41% compared to the state-of-the-art method, as shown in Table 1.
+The VIGOR results for SkyDiffusion use only a single street-view image.
+The results using Multi-to-One BEV generation with multiple street-view images can be found in the ablation study.
+The visual results in Fig.
+5 show that our method maintains consistent associative information such as crosswalks and lanes.
+While the non-associative information on building rooftops cannot be consistently generated, it still produces reasonable and realistic results.
+Method FID ($\downarrow$) SSIM ($\uparrow$) PSNR ($\uparrow$) LPIPS ($\downarrow$)
+CVACT
+Baseline 62.21 0.115 11.95 0.682
+Ours (BEV) 42.84 0.117 13.62 0.673
+Ours (C-BEV) 36.48 0.118 12.85 0.645
+VIGOR-Chicago
+Baseline 53.27 0.170 10.38 0.666
+Ours (BEV) 48.63 0.175 11.01 0.687
+Ours (C-BEV) 45.29 0.186 11.69 0.661
+Ours (C-BEV Multi) 31.90 0.205 11.15 0.651
+Table 2: Ablation study of the Curved-BEV module.
+"Baseline" represents directly using street-view image,"BEV" and "C-BEV" denotes using standard BEV or Curved-BEV transformation, and "Multi" stands for Multi-to-One.
+Method FID ($\downarrow$) SSIM ($\uparrow$) PSNR ($\uparrow$) LPIPS ($\downarrow$)
+Disaster Scene aerial image synthesis
+ControlNet 138.11 0.183 12.95 0.690
+Ours 96.16 0.195 13.07 0.685
+Low-altitude UAV image synthesis
+ControlNet 52.76 0.371 16.55 0.598
+Ours 32.53 0.452 16.86 0.390
+Historical HR satellite image synthesis
+ControlNet 36.48 0.117 12.85 0.665
+Ours 33.29 0.129 14.02 0.651
+Table 3: Quantitative comparison of methods on G2A-3 Dataset.
+5.3 Ablation Study
+We conducted ablation experiments on CVACT and VIGOR, with results shown in Table 2.
+Compared to directly using street-view images as input, the Curved-BEV method improves performance across multiple metrics by transforming street-view images into satellite views for domain alignment.
+This indicates that the Curved-BEV method aids in synthesizing more content-consistent satellite images.
+Furthermore, the Multi-to-one method further improves metrics compared to the one-to-one mapping, demonstrating its effectiveness in dense urban scenes.
+As shown in the visual results in Figure 6, using multiple street-view images for multi-to-one BEV mapping effectively addresses occlusions from tall objects like trees and buildings.
+This method effectively enhances the BEV sensing range in urban environments, facilitating the generation of large-scale satellite scenes.
+Additional ablation study results can be found in the supplementary materials.
+We will include additional experiments in the supplementary materials, such as comparisons between Curve-BEV and the standard BEV, as well as experiments on the hyperparameter $\lambda$, among others.
+5.4 Evaluation results on G2A-3 dataset
+As shown in Table 3, the tasks on the G2A-3 dataset present certain challenges; however, our method achieves significant performance improvements over the commonly used image-conditioned synthesis method, ControlNet.
+SkyDiffusion reduces the FID by an average of 25.81% and increases the SSIM by an average of 12.88%.
+ControlNet generates satellite images from street-view input, creating a significant domain gap.
+In contrast, our method uses Curve-BEV transformation to bridge the viewpoint domain gap, enabling more consistent satellite image generation.
+Disaster Scene Aerial image.
+In Figure 7(a), the key disaster-stricken regions, highlighted within the orange boxes, are distinctly visible in the generated satellite imagery.
+This provides a clear and comprehensive view of the disaster areas, with the potential to assist in subsequent assessment and in-depth analysis of the affected regions.
+Low-Altitude UAV Image.
+The example in Figure 7(b) demonstrates SkyDiffusion’s remarkable ability to synthesize low-altitude UAV images.
+Due to the lower altitude of UAV captures, there is greater overlap and closer domain similarity between scenes, resulting in superior synthesis quality.
+SkyDiffusion showcases outstanding aerial image synthesis capabilities, expanding its potential applications.
+Historical High-Resolution Satellite image.
+Figure 7(c) illustrates that between 2008 and 2014, the road underwent renovations, including the addition of crosswalks.
+SkyDiffusion successfully synthesizes realistic historical images from earlier periods, effectively addressing the absence of high-resolution Satellite imagery from those times.
+Additional visualizations on G2A-3 are available in the supplementary materials.
+Figure 7: Qualitative Visualization Results on the G2A-3 Dataset.
+6 Conclusion
+In this study, we introduce SkyDiffusion, a novel approach specifically designed for street images to satellite images cross-view synthesis.
+SkyDiffusion operates solely with street images as input, utilizing a BEV Paradigm and diffusion models to generate satellite images.
+SkyDiffusion achieves state-of-the-art performance in both content consistency and image realism on across multiple cross-view datasets, demonstrating its superior capabilities.
+Additionally, we introduce a cross-view synthesis dataset, Ground2Aerial-3, featuring aerial image synthesis tasks for multiple new scenes, providing practical value and inspiration for future cross-view synthesis research.
+7 Acknowledgement
+This work was supported in part by the Natural Science Foundation of Guangdong Province, China (Grant No.
+2025A1515010400), the National Natural Science Foundation of China (Grant No.
+42201358) and Shanghai Artificial Intelligence Laboratory.